### v2

# Courcelle’s Theorem for Lipschitz Continuity

Abstract
Lipschitz continuity of algorithms, introduced by Kumabe and Yoshida (FOCS’23), measures the stability of an algorithm against small input perturbations. Algorithms with small Lipschitz continuity are desirable, as they ensure reliable decision-making and reproducible scientific research. Several studies have proposed Lipschitz continuous algorithms for various combinatorial optimization problems, but these algorithms are problem-specific, requiring a separate design for each problem.
To address this issue, we provide the first algorithmic meta-theorem in the field of Lipschitz continuous algorithms. Our result can be seen as a Lipschitz continuous analogue of Courcelle’s theorem, which offers Lipschitz continuous algorithms for problems on bounded-treewidth graphs. Specifically, we consider the problem of finding a vertex set in a graph that maximizes or minimizes the total weight, subject to constraints expressed in monadic second-order logic ($\textsf{MSO}_{2}$). We show that for any $\varepsilon>0$, there exists a $(1\pm\varepsilon)$-approximation algorithm for the problem with a polylogarithmic Lipschitz constant on bounded treewidth graphs. On such graphs, our result outperforms most existing Lipschitz continuous algorithms in terms of approximability and/or Lipschitz continuity. Further, we provide similar results for problems on bounded-clique-width graphs subject to constraints expressed in $\textsf{MSO}_{1}$. Additionally, we construct a Lipschitz continuous version of Baker’s decomposition using our meta-theorem as a subroutine.
1 Introduction
Lipschitz continuity of algorithms, introduced by Kumabe and Yoshida [15], is a measure of an algorithm’s stability in response to errors or small perturbations in the input for weighted optimization problems. Roughly speaking, it is the maximum ratio of the (weighted) Hamming distance between the outputs of an algorithm for two different weights to the $\ell_{1}$ distance between those weights (see Section 2.4 for the precise definition). It is desirable for algorithms to have small Lipschitz constants, as large constants can undermine the reliability of decision-making and the reproducibility of scientific research.
Since its introduction, Lipschitz continuous algorithms have been proposed for various optimization problems [15, 17]. However, we need to design a different algorithm and analyze its Lipschitz continuity for each problem, which can be impractical. To address this limitation, we present an algorithmic meta-theorem for Lipschitz continuous algorithms, which can be seen as Lipschitz continuity analogue of celebrated Courcelle’s theorem [3].
To present our results, we first introduce some notation. Let $G=(V,E)$ be a graph of $n$ vertices with treewidth $\mathrm{tw}$, $\varphi(X)$ be an $\textsf{MSO}_{2}$ formula with a free vertex set variable $X$ (see Section 2.3 for details about logic), and $w\in\mathbb{R}_{\geq 0}^{V}$ be a weight vector. We consider the problem of finding a vertex subset $X\subseteq V$ such that $G\models\varphi(X)$ and $w(X)$ is maximized, which we call the $\textsf{MSO}_{2}$ maximization problem. We also consider the problem of finding $X$ such that $w(X)$ is minimized, which we call the $\textsf{MSO}_{2}$ minimization problem. Our main results are the following.
Theorem 1.
For any $\varepsilon\in(0,1]$, there is a $(1-\varepsilon)$-approximation algorithm for the $\textsf{MSO}_{2}$ maximization problem with Lipschitz constant $O\left((f(\mathrm{tw},|\varphi|)+\log\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n\right)$, where $f$ is some computable function. The time complexity is bounded by $O\left(f(\mathrm{tw},|\varphi|)n\right)$.
Theorem 2.
For any $\varepsilon\in(0,1]$, there is a $(1+\varepsilon)$-approximation algorithm for the $\textsf{MSO}_{2}$ minimization problem with Lipschitz constant $O\left((f(\mathrm{tw},|\varphi|)+\log\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n\right)$, where $f$ is some computable function. The time complexity is bounded by $O\left(f(\mathrm{tw},|\varphi|)n\right)$.
We note that the trivial upper bound on Lipschitz constant is $n$ [17]; therefore the bounds in the theorems above are significantly smaller for fixed $\mathrm{tw}$ and $\varphi$. We also remark that our meta-theorems yield randomized approximation algorithms. This is necessary since, for most problems, it is known that exact or deterministic algorithms cannot be Lipschitz continuous [15].
When the treewidth of the input graph is bounded by a constant, Theorems 1 and 2 provide algorithms that outperform existing algorithms in terms of approximability and/or Lipschitz continuity:
  • For the minimum weight vertex cover problem, Theorem 2 yields an algorithm with a better approximation ratio than the previous $2$-approximation algorithm with Lipschitz constant $4$ [17], at the polylogarithmic sacrifice of the Lipschitz constant.
  • For the minimum weight feedback vertex set problem, Theorem 2 outperforms the previous $O(\log n)$-approximation algorithm with Lipschitz constant $O(\sqrt{n}\log^{3/2}n)$ [17] in terms of both approximability and Lipschitz continuity.
  • For the maximum weight matching problem [^1] Although the output of this problem is an edge set rather than a vertex set, this problem can be expressed by an $\textsf{MSO}_{2}$ formula on a new graph $G^{\prime}=(V\cup E,\{(v,e)\colon e\text{ is incident to }v\})$, Theorem 1 yields an algorithm with a better approximation ratio than the previous $\left(\frac{1}{8}-\varepsilon\right)$-approximation algorithm with Lipschitz constant $O(\varepsilon^{-1})$ [15], at the polylogarithmic sacrifice of the Lipschitz constant.
  • For the shortest path problem, Theorem 2 slightly improves the Lipschitz continuity compared to the previous $(1-\varepsilon)$-approximation algorithm with Lipschitz constant $O(\varepsilon^{-1}\log^{3}n)$ [15], without losing approximability.
For a fixed $\varphi$, by explicitly specifying the transitions in the dynamic programming within the algorithm, we can provide more precise bounds on the function $f$ that appears in Theorems 1 and 2. In particular, considering the case that $\varphi$ is an $\textsf{MSO}_{2}$ formula representing the independent set constraint, we have the following.
Theorem 3.
For any $\varepsilon\in(0,1]$, there is a $(1-\varepsilon)$-approximation algorithm for the maximum weight independent set problem with Lipschitz constant $O\left((\mathrm{tw}+\log\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n\right)$. The time complexity is bounded by $2^{O(\mathrm{tw})}n$.
This result is surprising as the dependence of $\mathrm{tw}$ on the Lipschitz constant is subexponential and, even more remarkably, linear. We note that through similar arguments, analogous results hold for several other problems, such as the minimum weight vertex cover problem and the minimum weight dominating set problem.
We further demonstrate that Theorems 1 and 2 lead to Lipschitz continuous version of Baker’s technique [1]. As a representative example, we consider the maximum weight independent set problem on planar graphs, where a vertex subset $X$ is an independent set of a graph $G$ if no two vertices in $X$ are adjacent in $G$. We prove the following.
Theorem 4.
For any $\varepsilon\in(0,1]$, there is a $(1-\varepsilon)$-approximation algorithm for the maximum weight independent set problem on planar graphs with Lipschitz constant $O\left((\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n\right)$. The time complexity is bounded by $2^{O(\varepsilon^{-1})}n$.
Using similar algorithms, Lipschitz continuous PTASes can be obtained for many problems, including the minimum weight vertex cover problem and the minimum weight dominating set problem.
As a lower bound result, we prove that the Lipschitz constant of any $(1-\varepsilon)$-approximation algorithm for some $\textsf{MSO}_{2}$ maximization problem is large on general graphs, which justifies considering the $\textsf{MSO}_{2}$ maximization problem on a restricted graph class. Specifically, we consider the max ones problem [7, 10], where the instance is a 3CNF formula over a variable set $X$ and a weight function $w:X\to\mathbb{R}_{\geq 0}$, and the goal is to find a satisfying assignment $\sigma:X\to\{0,1\}$ that maximizes the weight $\sum_{x\in\sigma^{-1}(1)}w(x)$. It is easy to see that there exists a fixed $\textsf{MSO}_{2}$ formula $\varphi(X)$ such that the max ones problem can be reduced to a problem on bipartite graphs, where the task is to find a vertex set $X$ that maximizes the weight $\sum_{x\in X}w(x)$ subject to $\varphi$. We show the following.
Theorem 5.
There exist $\varepsilon,\delta>0$ such that any $(1-\varepsilon)$-approximation algorithm for the max ones problem has Lipschitz constant $\Omega(n^{\delta})$, where $n$ is the number of variables.
We note that it is possible to construct an algorithm for the max ones problem with a Lipschitz constant similar to that in Theorem 3, where $\mathrm{tw}$ denotes the treewidth of the “incidence graph” of the 3CNF instance (see Section 4.2 for details). Since $\mathrm{tw}$ is upper bounded by $n$ for any instance, Theorem 5 implies that the dependency of the Lipschitz constant on $\mathrm{tw}$ cannot be improved to $\mathrm{tw}^{o(1)}$.
We also prove analogous meta-theorems when parameterizing by clique-width. Let an $\textsf{MSO}_{1}$ maximization (resp., minimization) problem be the variant of an $\textsf{MSO}_{2}$ problem in which the formula $\varphi$ is restricted to $\textsf{MSO}_{1}$. Denote by $\mathrm{cw}$ the clique-width of the graph $G$. We then obtain the following.
Theorem 6.
For any $\varepsilon\in(0,1]$, there is a $(1-\varepsilon)$-approximation algorithm for the $\textsf{MSO}_{1}$ maximization problem with Lipschitz constant $O\left((f(\mathrm{cw},|\varphi|)+\log\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n\right)$, where $f$ is some computable function. The time complexity is bounded by $O\left(f(\mathrm{cw},|\varphi|)n\right)$.
Theorem 7.
For any $\varepsilon\in(0,1]$, there is a $(1+\varepsilon)$-approximation algorithm for the $\textsf{MSO}_{1}$ minimization problem with Lipschitz constant $O\left((f(\mathrm{cw},|\varphi|)+\log\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n\right)$, where $f$ is some computable function. The time complexity is bounded by $O\left(f(\mathrm{cw},|\varphi|)n\right)$.
We note that the functions $f$ in Theorems 6 and 7 are much larger than those in Theorems 1 and 2. In particular, the bounds for Theorems 3 and 4 do not follow from Theorem 6.
1.1 Technical Overview
Now we provide a technical overview of our framework. Since the arguments for clique-width are similar, we focus here on the treewidth results and omit the clique-width case.
For simplicity, here we consider the maximum weight independent set problem on a full binary tree (a rooted tree in which every vertex has exactly 0 or 2 children). This corresponds to the case where $\varphi(X)=\forall x\forall y((x\in X\land y\in X)\rightarrow\lnot\mathsf{adj}(x,y))$. Let $w\in\mathbb{R}_{\geq 0}^{V}$ be the weight vector.
If we do not care about Lipschitzness, this problem can exactly be solved by the following algorithm. For each vertex $v\in V$, define $\mathsf{DP}[v][0]$ to be an independent set in the subtree rooted at $v$ that does not include $v$ with the maximum weight. Similarly, define $\mathsf{DP}[v][1]$ to be an independent set in the subtree rooted at $v$ with the maximum weight. If $v$ is a leaf, then $\mathsf{DP}[v][0]=\emptyset$ and $\mathsf{DP}[v][1]=\{v\}$. Otherwise, let $u_{1}$ and $u_{2}$ be the two children of $v$. We have $\mathsf{DP}[v][0]=\mathsf{DP}[u_{1}][1]\cup\mathsf{DP}[u_{2}][1]$, and $\mathsf{DP}[v][1]$ is the one with the larger weight of $X_{v,0}:=\mathsf{DP}[u_{1}][0]\cup\mathsf{DP}[u_{2}][0]\cup\{v\}$ and $X_{v,1}:=\mathsf{DP}[u_{1}][1]\cup\mathsf{DP}[u_{2}][1]$. By performing this dynamic programming in a bottom-up manner, the problem can be solved exactly.
However, this algorithm is not Lipschitz continuous. This is because when we compute $\mathsf{DP}[v][1]$, which of $X_{v,0}$ or $X_{v,1}$ has a larger weight is affected by small changes in the weights. To address this issue, we use the exponential mechanism [8]. Specifically, for some constant $c>0$, we select $X_{v,i}$ with probability proportional to $\exp(c\cdot w(X_{v,i}))$. This approach makes the algorithm Lipschitz continuous, with only a slight sacrifice in the approximation ratio. Specifically, we can prove that by appropriately choosing $c$ for a given $\varepsilon^{\prime}\in(0,1]$, the increase in the Lipschitz constant can be bounded by $\tilde{O}(\varepsilon^{\prime-1})$ by reducing the approximation guarantee by a factor of $1-\varepsilon^{\prime}$ at each vertex where the exponential mechanism is applied.
While this approach makes the algorithm Lipschitz continuous, the Lipschitz constant is still too large when the height $h$ of the tree is large. Specifically, to achieve an approximation guarantee of $1-\varepsilon$, we need to set $\varepsilon^{\prime}=\frac{\varepsilon}{h}$, leading to a Lipschitz constant of $h\cdot\tilde{O}(\varepsilon^{\prime-1})=\tilde{O}(\varepsilon^{-1}h^{2})$. This is larger than the trivial bound $n$ when $h=O(n)$. We resolve this issue by using the fact that any tree has a tree decomposition of width 5 and height $O(\log n)$ [2]. By performing dynamic programming with the exponential mechanism on this tree decomposition, we can obtain $(1-\varepsilon)$-approximation algorithm with Lipschitz constant $\tilde{O}(\varepsilon^{-1})$. This argument can be naturally extended to the case where $G$ is a bounded treewidth graph. By following the proof of Courcelle’s Theorem [3, 6], we further extend this argument to the case where $\varphi$ is a general $\textsf{MSO}_{2}$ formula with a free vertex variable.
We prove Theorem 5 by leveraging a lower bound on a related notion of sensitivity for the maximum cut problem [9], where sensitivity measures the change in the output with respect to the Hamming distance under edge deletions [24]. Specifically, we reduce the maximum cut problem, where stability is defined with respect to edge deletions, to the max ones problem, where stability is measured with respect to weight changes.
1.2 Related Work
Lipschitz continuity of discrete algorithms is defined by Kumabe and Yoshida [15], and they provided Lipschitz continuous approximation algorithms for the minimum spanning tree, shortest path, and maximum weight matching problems. In a subsequent work [17], they further provided such algorithms for the minimum weight vertex cover, minimum weight set cover, and minimum weight feedback vertex set problems. In another work [16], they defined Lipschitz continuity for allocations in cooperative games and provided Lipschitz continuous allocation schemes for the matching game and the minimum spanning tree game.
A variant known as pointwise Lipschitz continuity has also been studied, which is defined using the unweighted Hamming distance instead of the weighted Hamming distance. Kumabe and Yoshida [15] defined pointwise Lipschitz continuity and provided pointwise Lipschitz continuous algorithms for the minimum spanning tree and maximum weight bipartite matching problems. Liu et al. [19] proposed the proximal gradient method as a general technique for solving LP relaxations stably. Using this, they provided pointwise Lipschitz continuous approximation algorithms for the minimum vertex $(S,T)$-cut, densest subgraph, maximum weight ($b$-)matching, and packing integer programming problems.
(Average) sensitivity, introduced by Varma and Yoshida [24] with preliminary ideas by Murai and Yoshida [20], is a notion similar to Lipschitz continuity. While Lipschitz continuity evaluates an algorithm’s stability against unit changes in the input weights, (average) sensitivity evaluates an algorithm’s stability against (random) deletions of elements from the input. Algorithms with small (average) sensitivity have been studied for several problems, such as maximum matching [24, 26], minimum cut [24], knapsack problem [14], Euclidean $k$-means [25], spectral clustering [22], and dynamic programming problems [13]. Recently, Fleming and Yoshida [9] constructed a PCP framework to prove the sensitivity lower bound for the constraint satisfaction problem.
1.3 Organization
The rest of this paper is organized as follows. In Section 2, we describe the necessary preliminaries on tree decomposition (Sections 2.1 and 2.2), logic (Section 2.3), and Lipschitz continuity (Section 2.4). In Section 3, we prove Theorem 1 by providing a Lipschitz continuous algorithm for the $\textsf{MSO}_{2}$ maximization problem. Since the algorithm and analysis for $\textsf{MSO}_{2}$ minimization are similar, we defer the proof of Theorem 2 to Appendix C. In Section 4, we give a more precise Lipschitzness analysis for specific formulas $\varphi$, which includes the proof of Theorem 3. In Section 5, we prove Theorem 4 by providing a Lipschitz continuous version of Baker’s technique. In Section 6, we discuss a lower bound on the Lipschitz constant for general graphs, which justifies focusing on graphs with bounded treewidth. In Section 7, we prove Theorem 6 by providing a Lipschitz continuous algorithm for the $\textsf{MSO}_{1}$ maximization problem, while we defer the proof of Theorem 7 to Appendix C.
2 Preliminaries
2.1 Tree Decomposition
Let $G$ be a graph with $n$ vertices. A pair $(\mathcal{B},\mathcal{T})$ consisting of a family $\mathcal{B}$ of subsets (called bags) of $V(G)$ and a rooted tree $\mathcal{T}$ whose vertex set is $\mathcal{B}$ is a (rooted) tree decomposition of $G$ if it satisfies the following three conditions.
  • $\bigcup_{B\in\mathcal{B}}B=V(G)$.
  • For each edge $e\in E(G)$, there is a bag $B\in\mathcal{B}$ such that $e\subseteq B$.
  • For each vertex $v\in V(G)$, the set of bags $\{B\in\mathcal{B}\colon v\in B\}$ induces connected subgraph in $T$.
For a tree decomposition $(\mathcal{B},\mathcal{T})$, we may refer to the root node, leaf nodes, and the height of $\mathcal{T}$ as those of $(\mathcal{B},\mathcal{T})$, respectively. Moreover, $(\mathcal{B},\mathcal{T})$ is binary if $\mathcal{T}$ is a binary tree. The width of a tree decomposition $(\mathcal{B},\mathcal{T})$ is the maximum size of a bag in $\mathcal{B}$ minus one. The treewidth of $G$ is the minimum possible width among all possible tree decompositions of $G$. It is known that the tree decomposition of $G$ of width at most $12k+$ 1 can be computed in $2^{O(k)}n$ time [11], where $k$ is the treewidth of $G$. Moreover, any tree decomposition of $G$ of width $k$ can be transformed into a binary tree decomposition of width at most $23k+$ 2 and height at most $O(\log n)$ [2]. Thus, we obtain the following lemma.
Lemma 8 ([2, 11]).
Let $G$ be a graph with $n$ vertices and $k$ be the treewidth of $G$. Then, a binary tree decomposition of $G$ with width $O(k)$ and height $O(\log n)$ can be computed in $2^{O(k)}n$ time.
2.2 HR-algebra
We introduce the notions of HR-algebra [^2] The acronym HR stands for hyperedge replacement [4]., which is one of the algebraic definitions of tree decomposition. A $k$-graph is a tuple $G=\langle V,E,\mathsf{src}\rangle$ of a set of vertices $V$, a set of edges $E$, and a $k$-vector $\mathsf{src}\in(V\cup\{\mathtt{nil}\})^{k}$. We write the $i$-th element of a vector $\mathsf{src}$ by $\mathsf{src}(i)$. If $\mathsf{src}(i)=u$, we say that $u$ is a $i$-source. We write the set $\{\mathsf{src}(i):i\in[k]\}\setminus\{\mathtt{nil}\}$ by $\mathsf{src}(G)$.
Let $G$ and $H$ be $k$-graphs. The parallel-composition of $G$ and $H$, denoted by $G\mathbin{//}H$, is the graph generated by the following procedure. First, create the disjoint union of $G$ and $H$, and identify the (non-$\mathtt{nil}$) vertices $\mathsf{src}_{G}(i)$ and $\mathsf{src}_{H}(i)$ for each $i\in[k]$. Finally, remove all the self-loops and multi-edges. Let $B$ be a non-empty subset of $[k]$. The source forgetting operation $\mathsf{fg}_{B}$ is the function that maps a $k$-graph $G$ to a $k$-graph $G^{\prime}$ such that $V_{G}=V_{G^{\prime}}$, $E_{G}=E_{G^{\prime}}$, and $\mathsf{src}_{G^{\prime}}(i)=\mathtt{nil}$ if $i\in B$, and $\mathsf{src}_{G^{\prime}}(i)=\mathsf{src}_{G}(i)$ otherwise.
Definition 9 (HR-algebra).
A term of a HR-algebra over $k$-graphs is either
  • a constant symbol $\mathtt{i}$, $\mathtt{ij}$, or $\emptyset$ denoting a $k$-graph $\langle\{v\},\emptyset,\mathsf{src}\rangle$ with $\mathsf{src}(i)=v$, a $k$-graph $\langle\{u,v\},\{\{u,v\}\},\mathsf{src}\rangle$ with $\mathsf{src}(i)=u$ and $\mathsf{src}(j)=v$, or the empty graph, respectively;
  • $t\mathbin{//}s$ for any terms $t$ and $s$;
  • $\mathsf{fg}_{B}(t)$ for any term $t$ and any $B\subseteq[k]$.
We may associate a term with the graphs represented by it. A term of an HR-algebra $t$ can be decomposed into the parse tree in the usual way, called the (rooted) HR-parse tree of $t$. The height of a HR-parse tree is the maximum distance from the root to a leaf. It is known that tree decompositions are equivalent to HR-parse trees in the following sense.
Proposition 10 (see e.g. [4]).
The treewidth of a graph is at most $k$ if and only if the graph can be denoted by a term of a HR-algebra over $(k+1)$-graphs.
Moreover, binary tree decomposition can be transformed into an HR-parse tree with approximately the same height.
Proposition 11 ($\star$).
Given a rooted binary tree decomposition $(\mathcal{B},\mathcal{T})$ of $G$ with width $k$ and height $h$, we can compute a HR-parse tree over $(k+1)$-graph denoting $G$ with height $O(\log k+h)$.
The proof of Proposition 11 is essentially the same as the proof of Proposition 10 but we give in Appendix A.
2.3 Monadic Second-Order Logic
We provide a slightly less formal definition of monadic second-order (MSO) logic over $k$-graphs for accessibility (see, e.g., [18] for more formal definition). An atomic formula is a formula of the form $s=t$, $\mathsf{adj}(s,t)$, $t\in X$, and Boolean constants $\mathtt{True}$ and $\mathtt{False}$. Here, $\mathsf{adj}(s,t)$ is the predicate that represents whether $s$ and $t$ are adjacent. A monadic second-order formula over $k$-graphs is a formula built from atomic formulas, the usual Boolean connectives ,,¬$\land$,$\lor$,$\lnot$,$\to$ , ∨ , ¬ , →, first-order quantifications $\exists x$ and $\forall x$ (quantifications over vertices), and second-order quantifications $\exists X$ and $\forall X$ (quantification over sets of vertices). The notation $\varphi(X)$ indicates that $\varphi$ has exactly one free variable $X$. The quantifier height (or quantifier rank) of a formula $\varphi$, denoted by $|\varphi|$, is the maximum depth of quantifier nesting in $\varphi$ (see e.g. [18] for a formal definition). The counting monadic second-order (CMSO) logic is an expansion of MSO logic that have additional predicates $\mathsf{Card}_{m,r}(X)$ for a second-order variable $X$, meaning $|X|\equiv m\pmod{r}$. A $\textsf{C}_{r}\textsf{MSO}^{q}$-formula is a CMSO formula $\varphi$ such that the quantifier height of $\varphi$ is at most $q$ and $r^{\prime}\leq r$ holds for any predicate $\mathsf{Card}_{m,r^{\prime}}$ in $\varphi$. For a logical formula $\varphi$ and a graph $G$, we write $G\models\varphi$ when graph $G$ satisfies the property expressed by $\varphi$.
There are two variants of MSO logic: $\textsf{MSO}_{1}$ (or simply MSO), which is described above, allows quantification over vertices and sets of vertices only, and $\textsf{MSO}_{2}$ which allows quantification over edges and sets of edges as well. It is well known that, for any graph $G$ and $\textsf{MSO}_{2}$-formula $\varphi$, there exists an $\textsf{MSO}_{1}$ formula $\varphi^{\prime}$ such that $G\models\varphi\iff G^{\prime}\models\varphi^{\prime}$, where $G^{\prime}$ is the incidence graph, obtained by subdividing each edge of $G$ (with two colors meaning the original and the subdivision vertices) (see e.g. [12]). It is easy to see that the treewidth of the incidence graph is at most the treewidth of the original graph. Thus, since this paper focuses on graphs with bounded treewidth, we mainly consider $\textsf{MSO}_{1}$. However, all the theorems are held for $\textsf{MSO}_{2}$ and graphs of bounded treewidth.
We introduce some more notations. For a set $V$, two families $\mathcal{A}\subseteq 2^{V}$ and $\mathcal{B}\subseteq 2^{V}$ are separated if $A\cap B=\varnothing$ for all $A\in\mathcal{A}$, and $B\in\mathcal{B}$. We write $A\cup B$ by $A\uplus B$ if $A\cap B=\varnothing$, and $\mathcal{A}\boxtimes\mathcal{B}\coloneqq\{A\cup B:A\in\mathcal{A},B\in\mathcal{B}\}$ if $\mathcal{A}$ and $\mathcal{B}$ are separated. For a formula $\varphi(X)$ and a graph $G$, the set of vertex sets satisfying $\varphi$ is denoted by $\mathsf{sat}(G,\varphi)$, that is, $\mathsf{sat}(G,\varphi)=\{A:G\models\varphi(A)\}$.
The following theorem is a key of the MSO model-checking algorithm [3, 4].
Theorem 12 ([4]).
Let $r,q,k$ be positive integers. Let $\varphi({X})$ be a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula over $k$-graphs with a second-order free variable ${X}$. Then, the following hold.
  1. 1.
    For any $B\subseteq[k]$, there exists a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula $\psi({X})$ such that for any $k$-graph $G$, we have
    $\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi)=\mathsf{sat}(G,\psi).$
  2. 2.
    There exists a family of tuples $\{\theta_{i}({X}),\psi_{i}({X})\}_{i\in[p]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with a free variable ${X}$ such that, for any $k$-graphs $G$ and $H$,
    $\mathsf{sat}(G\mathbin{//}H,\varphi)=\biguplus_{i\in[p]}\{S\cup P:S\in\mathsf{sat}(G,\theta_{i}),P\in\mathsf{sat}(H,\psi_{i})\}.$
It is known that, for any $r,q,k$, up to logical equivalence, there are only finitely many different $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas $\varphi(X)$ over $k$-graphs (see e.g., [18]). Thus, we can obtain a linear-time algorithm, based on dynamic programming over trees, for $\textsf{C}_{r}\textsf{MSO}^{q}$ model checking over bounded treewidth. Courcelle and Mosbah [6] adjusted Theorem 12 to address optimization, counting, and other problems. We use a slightly modified version of the theorem of Courcelle and Mosbah.
Let $\mathsf{nosrc}(X)\equiv\bigwedge_{i\in[k]}(\mathsf{src}(i)\not\in X)$, and $\varphi\mathord{{\restriction}_{S}}(X)$ denote $\varphi(X\cup S)\land\mathsf{nosrc}(X)$ for a set $S\subseteq\mathsf{src}(G)$.
Observation 13.
Let $G$ be a $k$-graph, and $\varphi(X)$ be a formula with a free set variable $X$. Then, we have $\mathsf{sat}(G,\varphi)=\biguplus_{S\subseteq\mathsf{src}(G)}\mathsf{sat}(G,\varphi\mathord{{\restriction}_{S}})\boxtimes\{S\}$.
The proof of Corollary 14 is given in Appendix A
Corollary 14 ($\star$).
Let $r,q,k$ be positive integers. Let $\varphi({X})$ be a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula over $k$-graphs with a second-order free variables ${X}$. Then, for all $S\subseteq\mathsf{src}(G)$, the following hold.
  1. 1.
    For any $B\subseteq[k]$, there exists a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula $\psi({X})$ such that for any $k$-graph $G$, we have
    $\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi\mathord{{\restriction}_{S}})=\biguplus_{S^{\prime}\subseteq B}\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\}.$
  2. 2.
    There exists a family of tuples $\{\theta_{i}\mathord{{\restriction}_{S_{i}}}({X}),\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}({X})\}_{i\in[p]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with a free variable ${X}$ such that, for any $k$-graphs $G$ and $H$,
    $\mathsf{sat}(G\mathbin{//}H,\varphi\mathord{{\restriction}_{S}})=\biguplus_{i\in[p]}\mathsf{sat}(G,\theta_{i}\mathord{{\restriction}_{S_{i}}})\boxtimes\mathsf{sat}(H,\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}).$
    Moreover, $S_{i}\cup S^{\prime}_{i}=S$ for all $i\in[p]$.
Then, we can design an algorithm for an (C)MSO maximization (minimization) problem over graphs of bounded treewidth. For simplicity, we assume that the given graph $G$ has no sources, that is, $\mathsf{src}(G)=\varnothing$. Let $t$ be a term denoting $G_{t}$ and $\varphi(X)$ be the (C)MSO-formula describing the constraint of the problem. We recursively compute the value $\mathsf{opt}(t^{\prime},\varphi^{\prime}\mathord{{\restriction}_{S}})=\max\{w(A):G\models\varphi^{\prime}\mathord{{\restriction}_{S}}(A)\}$ for any subterm $t^{\prime}$ of $t$ and MSO-formula $\varphi^{\prime}\mathord{{\restriction}_{S}}$ as follows, where $w\in\mathbb{R}_{\geq 0}^{V}$ is the given weight vector and $w(X)=\sum_{x\in X}w_{x}$ for any $X\subseteq V(G)$. If $t^{\prime}$ is of the form $\mathsf{fg}_{B}(t^{\prime\prime})$, then $\mathsf{opt}(\mathsf{fg}_{B}(t^{\prime\prime}),\varphi^{\prime}\mathord{{\restriction}_{S}})=\max_{S^{\prime}\subseteq B}\{\mathsf{opt}(t^{\prime\prime},\psi\mathord{{\restriction}_{S\cup S^{\prime}}})+w(S^{\prime})\}$, where, $\psi$ is the formula obtained from Corollary 14. If $t^{\prime}$ is of the form $t_{1}\mathbin{//}t_{2}$, then $\mathsf{opt}(t^{\prime},\varphi^{\prime}\mathord{{\restriction}_{S}})=\max_{i\in[p]}\{\mathsf{opt}(t_{1},\theta\mathord{{\restriction}_{S_{i}}})+\mathsf{opt}(t_{2},\psi\mathord{{\restriction}_{S^{\prime}_{i}}})\}$, where, $\{\theta_{i}\mathord{{\restriction}_{S_{i}}}({X}),\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}({X})\}_{i\in[p]}$ are the formulas obtained from Corollary 14. Then, $\mathsf{opt}(t,\varphi\mathord{{\restriction}_{\varnothing}})$ is the maximum value for the problem since $t$ has no sources. In the following, for simplicity, we may omit the $\mathord{{\restriction}_{S}}$ notation and consider only formulas $\varphi$ such that $\mathsf{sat}(G,\varphi)$ does not contain any sources.
In the above setting, the graphs are considered without any special vertex sets (called as colors) or special vertices (called as labels). However, for $k$-graphs with a constant number of colors and/or labels, an appropriate HR-algebra and logics can be defined, and the similar results for Corollary 14 hold (see e.g. [4, 18]).
2.4 Lipschitz Continuity
We formally define Lipschitz continuity of algorithms. As this is sufficient for this work, we only consider algorithms for vertex-weighted graph problems. The definition can be naturally extended to other settings, such as edge-weighted problems. Let $G=(V,E)$ be a graph. For vertex sets $X,X^{\prime}\subseteq V$ and weight vectors $w,w^{\prime}\in\mathbb{R}_{\geq 0}^{V}$, we define the weighted Hamming distance between $(X,w)$ and $(X^{\prime},w^{\prime})$ by
$d_{\mathrm{w}}((X,w),(X^{\prime},w^{\prime})):=\left\|\sum_{v\in X}\bm{1}_{v}w_{v}-\sum_{v\in X^{\prime}}\bm{1}_{v}w^{\prime}_{v}\right\|_{1}=\sum_{v\in X\cap X^{\prime}}|w_{v}-w^{\prime}_{v}|+\sum_{v\in X\setminus X^{\prime}}w_{v}+\sum_{v\in X^{\prime}\setminus X}w^{\prime}_{v},$
where $\bm{1}_{v}\in\{0,1\}^{V}$ denotes the characteristic vector of $v$, that is, the vector $\bm{1}_{v,u}=1$ holds if and only if $u=v$. For two probability distributions $\mathcal{X}$ and $\mathcal{X}^{\prime}$ over subsets of $V$, we define
$\mathsf{EM}\left((\mathcal{X},w),(\mathcal{X}^{\prime},w^{\prime})\right):=\inf_{\mathcal{D}}\mathbb{E}_{(X,X^{\prime})\sim\mathcal{D}}\left[d_{\mathrm{w}}((X,w),(X^{\prime},w^{\prime}))\right],$
where the minimum is taken over couplings of $\mathcal{X}$ and $\mathcal{X}^{\prime}$, that is, distributions over pairs of sets such that its marginal distributions on the first and second coordinates are equal to $\mathcal{X}$ and $\mathcal{X}^{\prime}$, respectively. Consider an algorithm $\mathcal{A}$, that takes a graph $G=(V,E)$ and a weight vector $w\in\mathbb{R}_{\geq 0}^{V}$ as an input and outputs a vertex subset $X\subseteq V$. We denote the output distribution of $\mathcal{A}$ for weight $w$ as $\mathcal{A}(w)$. The Lipschitz constant of $\mathcal{A}$ is defined by
$\sup_{\begin{subarray}{c}w,w^{\prime}\in\mathbb{R}_{\geq 0},\\ w\neq w^{\prime}\end{subarray}}\frac{\mathsf{EM}\left((\mathcal{A}(w),w),(\mathcal{A}(w^{\prime}),w^{\prime})\right)}{\|w-w^{\prime}\|_{1}}.$
For two random variables $\bm{X}$ and $\bm{X}^{\prime}$, the total variation distance between them is given as:
$\mathsf{TV}\left(\bm{X},\bm{X}^{\prime}\right):=\inf_{\mathcal{D}}\Pr_{(X,X^{\prime})\sim\mathcal{D}}\left[X\neq X^{\prime}\right],$
where the minimum is taken over couplings between $\bm{X}$ and $\bm{X}^{\prime}$, that is, distributions over pairs such that its marginal distributions on the first and the second coordinates are equal to $\bm{X}$ and $\bm{X}^{\prime}$, respectively. For an element $u\in V$, we use $\bm{1}_{u}\in\mathbb{R}_{\geq 0}^{V}$ to denote the characteristic vector of $u$, that is, $\bm{1}_{u}(u)=1$ and $\bm{1}_{u}(v)=0$ for $v\in V\setminus\{u\}$. The following lemma indicates that, to bound the Lipschitz constant, it suffices to consider pairs of weight vectors that differ by one coordinate.
Lemma 15 ([15]).
Suppose that there exist some $c>0$ and $L>0$ such that
$\mathsf{EM}\left((\mathcal{A}(G,w),w),(\mathcal{A}(G,w+\delta\mathbf{1}_{u}),w+\delta\mathbf{1}_{u})\right)\leq\delta L$
holds for all $w\in\mathbb{R}_{\geq 0}^{V}$, $u\in V$ and $0<\delta\leq c$. Then, $\mathcal{A}$ is $L$-Lipschitz.
In our algorithm, we use the following procedures softmax and softmin. These procedures are derived from the exponential mechanism in the literature of differential privacy [8] and frequently appear in the literature on Lipschitz continuity [15, 17]. Here, we organize them into a more convenient form for our use. Let $p\in\mathbb{Z}_{\geq 1}$, $x_{1},\dots,x_{p}\in\mathbb{R}_{\geq 0}$, and $\epsilon\in(0,1]$. The softmax of $x_{1},\dots,x_{p}$ is taken as follows. If $\max_{i\in[p]}x_{i}=0$, we set $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}$ to be an arbitrary index $i$ with $x_{i}=0$ and $\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}=0$. Assume otherwise. First, we sample $\bm{c}$ uniformly from $\left[\frac{2\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}},\frac{4\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}}\right]$. Let $\bm{i}^{*}$ be a probability distribution over $[p]$ such that
$\Pr\left[\bm{i}^{*}=i\right]=\frac{\exp(\bm{c}x_{i})}{\sum_{i^{\prime}\in[p]}\exp(\bm{c}x_{i^{\prime}})}.$
holds for all $i\in[p]$. Then, we define $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}:=\bm{i}^{*}$ and $\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}:=x_{\bm{i}^{*}}$. We have the following. The proofs of Lemmas 16 and 17 are given in Appendix B.
Lemma 16.
We have $\mathbb{E}\left[\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}\right]\geq(1-\epsilon)\max_{i\in[p]}x_{i}$.
Lemma 17.
Let $\delta>0$ and assume $\delta\leq\frac{\max_{i\in[p]}x_{i}}{4\epsilon^{-1}\log(2p\epsilon^{-1})}$. Let $x_{1}^{\prime},\dots,x_{p}^{\prime}\in\mathbb{R}_{\geq 0}$ be numbers such that $x_{i}\leq x_{i}^{\prime}\leq x_{i}+\delta$ holds for all $i\in[p]$. Then, $\mathsf{TV}\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}$.
Similarly, the softmin of $x_{1},\dots,x_{p}$ is taken as follows. If $\min_{i\in[p]}x_{i}=0$, we set $\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}$ be an arbitrary index $i$ with $x_{i}=0$ and $\mathsf{softmin}^{\epsilon}_{i\in[p]}x_{i}=0$. Assume otherwise. First, we sample $\bm{c}$ uniformly from $\left[\frac{2\epsilon^{-1}\log(2p\epsilon^{-1})}{\min_{i\in[p]}x_{i}},\frac{4\epsilon^{-1}\log(2p\epsilon^{-1})}{\min_{i\in[p]}x_{i}}\right]$. Let $\bm{i}^{*}$ be a probability distribution over $[p]$ such that
$\Pr\left[\bm{i}^{*}=i\right]=\frac{\exp(-\bm{c}x_{i})}{\sum_{i^{\prime}\in[p]}\exp(-\bm{c}x_{i^{\prime}})}.$
holds for all $i\in[p]$. Then, we define $\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}:=\bm{i}^{*}$ and $\mathsf{softmin}^{\epsilon}_{i\in[p]}x_{i}:=x_{\bm{i}^{*}}$. We have the following. The proofs of Lemmas 18 and 19 are given in Appendix B.
Lemma 18.
We have $\mathbb{E}\left[\mathsf{softmin}^{\epsilon}_{i\in[p]}x_{i}\right]\leq(1+\epsilon)\min_{i\in[p]}x_{i}$.
Lemma 19.
Let $\delta>0$ and assume $\delta\leq\frac{\min_{i\in[p]}x_{i}}{4\epsilon^{-1}\log(2p\epsilon^{-1})}$. Let $x_{1}^{\prime},\dots,x_{p}^{\prime}\in\mathbb{R}_{\geq 0}$ be numbers such that $x_{i}\leq x_{i}^{\prime}\leq x_{i}+\delta$ holds for all $i\in[p]$. Then, $\mathsf{TV}\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}$.
3 Lipschitz Continuous Algorithms for $\textsf{MSO}_{2}$ Optimization Problems
In this section, we prove Theorems 1, 2 and 3. Since the proofs are similar to that for Theorem 1, we provide most of the proof of Theorem 2 in Appendix C. In Section 3.1, we define the notations and give a brief overview of the algorithm. Section 3.2 handles the base case, where the graph consists of a single vertex. Sections 3.3 and 3.4 analyze the impact of the parallel composition and the forget operations, respectively, on approximability and the Lipschitz constant. In Section 3.5, we put everything together to prove Theorem 1. Finally, in Section 4.1, we provide a more refined analysis for the special case of the maximum weight independent set problem to prove Theorem 3.
3.1 Definitions
For a graph $G$, a weight vector $w$, and an $\textsf{MSO}_{2}$ formula $\varphi$, we define
$\mathsf{opt}_{w}[G,\varphi]=\max_{S\subseteq V(G),G\models\varphi(S)}w(S).$ (1)
for maximization problems and
$\mathsf{opt}_{w}[G,\varphi]=\min_{S\subseteq V(G),G\models\varphi(S)}w(S).$ (2)
for minimization problems. If $\mathsf{sat}(G,\varphi)=\emptyset$, we define $\mathsf{opt}_{w}[G,\varphi]=\mathtt{nil}$. For graphs $G$ and $\textsf{MSO}_{2}$ formulas $\varphi$ with $\mathsf{sat}(G,\varphi)\neq\emptyset$, our algorithm recursively computes the vertex set $\mathsf{DP}_{w}[G,\varphi]$ that approximately achieves the maximum in Equation (1) or minimum in Equation (2). When it is clear from the context, we omit the subscript $w$ and simply write $\mathsf{opt}[G,\varphi]$ and $\mathsf{DP}[G,\varphi]$. We perform the dynamic programming algorithm using the formulas defined in Corollary 14 over the parse tree of a term obtained from Lemma 8 and Proposition 11. Specifically, at every stage of the algorithm, it is guaranteed that each $X\in\mathsf{sat}(G,\varphi)$ satisfies $X\subseteq V(G)\setminus\mathsf{src}(G)$. Furthermore, our algorithm is randomized. Thus, $\mathsf{DP}[G,\varphi]$ can be considered as a probability distribution over vertex sets in $\mathsf{sat}(G,\varphi)$.
To bound the approximation ratio, for a weight vector $w\in\mathbb{R}_{\geq 0}^{V}$, we bound the ratio between $\mathbb{E}\left[w(\mathsf{DP}[G,\varphi])\right]$ and $\mathsf{opt}[G,\varphi]$. To bound the Lipschitz constant, for a weight vector $w\in\mathbb{R}_{\geq 0}^{V}$, $u\in V$, and $\delta>0$, we bound $\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi]\right)$. From now on, we will concentrate on maximizing problems. The algorithm for minimization problems is similar to that for maximization problems, while the detail of the analysis is slightly differ. We discuss the minimization version in Appendix C.
3.2 Base Case
Here we consider the base case. Let $G$ be a graph with a single vertex $v$ and no edges, where $v$ is a source of $G$, and $\varphi$ be an $\textsf{MSO}_{2}$ formula such that $\mathsf{sat}(G,\varphi)$ is nonempty and contains no set containing $v$. In particular, we have $\mathsf{sat}(G,\varphi)=\{\emptyset\}$. We set $\mathsf{DP}[G,\varphi]:=\emptyset$. Since $w(\emptyset)=0$, it is clear that $w\left(\mathsf{DP}[G,\varphi]\right)=\mathsf{opt}[G,\varphi]=0$. Furthermore, it is obvious that
$\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi]\right)=\mathsf{EM}(\emptyset,\emptyset)=0.$
3.3 Parallel Composition
Next, we consider the parallel composition. Let
$\mathsf{sat}(G\mathbin{//}H,\varphi)=\biguplus_{i\in[p]}\mathsf{sat}(G,\theta_{i})\boxtimes\mathsf{sat}(H,\psi_{i}).$ (3)
Assume we have already computed $\mathsf{DP}[G,\theta_{i}]$ and $\mathsf{DP}[H,\psi_{i}]$ for each $i\in[p]$. We compute $\mathsf{DP}[G\mathbin{//}H,\varphi]$. For each $i\in[p]$, we denote $\mathsf{opt}_{i}:=\mathsf{opt}[G,\theta_{i}]+\mathsf{opt}[H,\psi_{i}]$. By definition, we have $\max_{i\in[p]}\mathsf{opt}_{i}=\mathsf{opt}[G\mathbin{//}H,\varphi]$. Then, we take $\bm{i}^{*}=\mathsf{softargmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}$ and define $\mathsf{DP}[G\mathbin{//}H,\varphi]:=\mathsf{DP}[G,\theta_{\bm{i}^{*}}]\cup\mathsf{DP}[H,\psi_{\bm{i}^{*}}]$. First we analyze the approximation ratio.
Lemma 20.
Let $0<\alpha\leq 1$ and suppose $\mathbb{E}\left[w(\mathsf{DP}[G,\theta_{i}])\right]\geq\alpha\mathsf{opt}[G,\theta_{i}]$ and $\mathbb{E}\left[w(\mathsf{DP}[H,\psi_{i}])\right]\geq\alpha\mathsf{opt}[H,\psi_{i}]$ hold for all $i\in[p]$. Then, we have $\mathbb{E}\left[w\left(\mathsf{DP}[G\mathbin{//}H,\varphi]\right)\right]\geq(1-\varepsilon)\alpha\mathsf{opt}[G\mathbin{//}H,\varphi]$.
Proof.
We have
$\mathbb{E}\left[w(\mathsf{DP}[G\mathbin{//}H,\varphi])\right]=\mathbb{E}\left[w(\mathsf{DP}[G,\theta_{\bm{i}^{*}}]\cup\mathsf{DP}[H,\psi_{\bm{i}^{*}}])\right]\geq\alpha\mathbb{E}\left[\mathsf{opt}[G,\theta_{\bm{i}^{*}}]+\mathsf{opt}[H,\psi_{\bm{i}^{*}}]\right]$
$=\alpha\mathbb{E}\left[\mathsf{softmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}\right]\geq(1-\varepsilon)\alpha\max_{i\in[p]}\mathsf{opt}_{i}=(1-\varepsilon)\alpha\mathsf{opt}[G\mathbin{//}H,\varphi].\qed$
Now we analyze the Lipschitz constant. We denote the variable $\bm{i}^{*}$ corresponding to the weight $w$ and $w+\delta\bm{1}_{u}$ by $\bm{i}^{*}_{w}$ and $\bm{i}^{*}_{w+\delta\bm{1}_{u}}$, respectively. We note that $\mathsf{DP}_{w}[G,\theta_{i}]$ and $\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\theta_{i}]$ are the same as a distribution unless $u\in V(G)\setminus\mathsf{src}_{G}$. The same also holds for $H$. Since $V(G)\setminus\mathsf{src}_{G}$ and $V(H)\setminus\mathsf{src}_{H}$ are disjoint, without loss of generality, we can assume the $\mathsf{DP}_{w}[H,\psi_{i}]$ and $\mathsf{DP}_{w+\delta\bm{1}_{u}}[H,\psi_{i}]$ are the same as a distribution. We have the following.
Lemma 21.
Let $\beta\in\mathbb{R}_{\geq 0}$ and suppose $\mathsf{EM}\left(\mathsf{DP}_{w}[G,\theta_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\theta_{i}]\right)\leq\beta$ holds for all $i\in[p]$. Then, we have $\mathsf{EM}\left(\mathsf{DP}_{w}[G\mathbin{//}H,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)\leq 30\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta$.
Proof.
We have
$\mathsf{EM}\left(\mathsf{DP}_{w}[G\mathbin{//}H,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)$
$\leq\mathsf{TV}\left(\bm{i}^{*}_{w},\bm{i}^{*}_{w+\delta\bm{1}_{u}}\right)\left(\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)+\max_{i\in[p]}\left(\mathsf{EM}\left(\mathsf{DP}_{w}[G,\theta_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\theta_{i}]\right)\right)$
$\leq\frac{10\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta}{\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]}\left(\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)+\beta\leq 30\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta,$
where the last inequality is from $\varepsilon\leq 1$ and $\delta\leq\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]$. ∎
3.4 Forget
Here we consider forget operation. Let $B\subseteq\mathsf{src}_{G}$ and
$\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi)=\biguplus_{i\in[p]}\mathsf{sat}(G,\varphi_{i})\boxtimes\{S_{i}\}.$ (4)
Assume we have already computed $\mathsf{DP}[G,\varphi_{i}]$ for each $i\in[p]$. We compute $\mathsf{DP}[\mathsf{fg}_{B}(G),\varphi]$.
We first sample $\bm{c}\in\mathbb{R}_{>0}$ uniformly from $\left[\frac{2\log(2p\varepsilon^{-1})}{\varepsilon\mathsf{opt}[G,\varphi,S]},\frac{4\log(2p\varepsilon^{-1})}{\varepsilon\mathsf{opt}[G,\varphi,S]}\right]$. For each $i\in[p]$, we denote $\mathsf{opt}_{i}=\mathsf{opt}[G,\varphi_{i}]+w(S_{i})$. By definition, we have $\max_{i\in[p]}\mathsf{opt}_{i}=\mathsf{opt}[\mathsf{fg}_{B}(G),\varphi]$. Then, we take $\bm{i}^{*}=\mathsf{softargmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}$ and define $\mathsf{DP}[\mathsf{fg}_{B}(G),\varphi]:=\mathsf{DP}[G,\varphi_{\bm{i}^{*}}]\cup S_{\bm{i}^{*}}$. First we analyze the approximation ratio.
Lemma 22.
Let $0<\alpha\leq 1$ and suppose $\mathbb{E}\left[w(\mathsf{DP}[G,\varphi_{i}])\right]\geq\alpha\mathsf{opt}[G,\varphi_{i}]$ holds for all $i\in[p]$. Then, we have $\mathsf{DP}[\mathsf{fg}_{B}(G),\varphi]\geq(1-\varepsilon)\alpha\mathsf{opt}[\mathsf{fg}_{B}(S),\varphi]$.
Proof.
We have
$\mathbb{E}\left[w(\mathsf{DP}[G,\varphi])\right]$ $=\mathbb{E}\left[w(\mathsf{DP}[G,\varphi_{\bm{i}^{*}}]\cup S_{\bm{i}^{*}})\right]\geq\mathbb{E}\left[\alpha\mathsf{opt}[G,\varphi_{\bm{i}^{*}}]+w(S_{\bm{i}^{*}})\right]\geq\alpha\mathbb{E}\left[\mathsf{opt}[G,\varphi_{\bm{i}^{*}}]+w(S_{\bm{i}^{*}})\right]$
$=\alpha\mathbb{E}\left[\mathsf{softmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}\right]\geq(1-\varepsilon)\alpha\max_{i\in[p]}\mathsf{opt}_{i}=(1-\varepsilon)\alpha\mathsf{opt}[\mathsf{fg}_{B}(G),\varphi].\qed$
Now we analyze the Lipschitz constant.
Lemma 23.
Let $\beta\in\mathbb{R}_{\geq 0}$ and suppose $\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi_{i}]\right)\leq\beta$ holds for all $i\in[p]$. Then, we have $\mathsf{EM}\left(\mathsf{DP}_{w}[\mathsf{fg}_{B}(G),\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)\leq 31\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta$.
Proof.
We have
$\mathsf{EM}$ $\left(\mathsf{DP}_{w}[\mathsf{fg}_{B}(G),\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)$
$\leq\mathsf{TV}\left(\bm{i}^{*}_{w},\bm{i}^{*}_{w+\delta\bm{1}_{u}}\right)\left(\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)$
$\quad+\max_{i\in[p]}\left(\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi_{i}]\cup S_{i},\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi_{i}]\cup S_{i}\right)\right)$
$\leq\frac{10\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta}{\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi]}\left(\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)$
$\quad+\max_{i\in[p]}\left(\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi_{i}]\right)\right)+\max_{i\in[p]}\left[d_{\mathrm{w}}((S_{i},w),(S_{i},w+\delta\bm{1}_{u}))\right]$
$\leq 30\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta+\delta\leq 31\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta.\qed$
3.5 Putting Together
Let $G=(V,E)$ be a graph with treewidth $k$. From Lemma 8 and Proposition 11, we can compute a HR-parse tree $t$ over $(k+1)$-graph denoting $G$ with height $O(\log k+\log n)\leq O(\log n)$. We perform the dynamic programming algorithm on this parse tree. We have the following.
Lemma 24.
Let $\varepsilon\in(0,1]$ and $h$ be the height of $t$. Our algorithm outputs a solution $X\in\mathsf{sat}(G,\varphi)$ such that $\mathbb{E}[w(X)]\geq(1-h\varepsilon)\mathsf{opt}[G,\varphi]$. The Lipschitz constant is bounded by $31h\varepsilon^{-1}\log(2p_{\max}\varepsilon^{-1})$, where $p_{\max}$ is the maximum of $p$ among all update formulas (3) or (4) that the algorithm uses.
Proof.
The approximability bound is obtained by repeatedly applying Lemmas 20 and 22 and $(1-\varepsilon)^{h}\geq 1-h\varepsilon$. The Lipschitzness bound is obtained by repeatedly applying Lemmas 21 and 23. ∎
Since $p_{\max}$ is bounded by a function of $k$ and $|\varphi|$, we have the following.
Proof of Theorem 1.
The claim follows by substituting $\varepsilon$, $h$, and $k$ in Lemma 24 with $\frac{\varepsilon}{\Theta(\log n)}$, $O(\log n)$, and $23k+$ 2, respectively. ∎
4 Special Cases
4.1 Independent Set
From the proof of Theorem 1, $f(k,\varphi)$ in Theorem 1 is bounded by $\log(2p_{\max})$. Particularly, if $p_{\max}$ is bounded by $2^{O(k)}$, the Lipschitz constant depends linearly on $k$. Here, we prove Theorem 3 for the maximum weight independent set problem by showing that $p_{\max}\leq 2^{O(k)}$. Similar arguments can be applied to several other problems, such as the minimum weight vertex cover problem and the minimum weight dominating set problem. For a $k$-graph $G$ and $S\subseteq\mathsf{src}(G)$, let
$\varphi_{S}(X)\equiv\forall x(x\in\mathsf{src}(G)\rightarrow\lnot x\in X)\land\forall x\forall y(((x\in X\lor x\in S)\land(y\in X\lor y\in S))\rightarrow\lnot\mathsf{adj}(x,y)).$
In words, $\mathsf{sat}(G,\varphi_{S})$ is the family of subsets $X$ of $V$ such that $X$ is disjoint from $\mathsf{src}(G)$ and $X\cup S$ is an independent set of $G$. Then, we have
$\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi_{S})=\biguplus_{S^{\prime}\subseteq B}\mathsf{sat}(G,\varphi_{S\cup S^{\prime}})\boxtimes\{S^{\prime}\},\quad\mathsf{sat}(G\mathbin{//}H,\varphi_{S})=\mathsf{sat}(G,\varphi_{S})\boxtimes\mathsf{sat}(H,\varphi_{S}).$
Particularly, we have $p_{\max}\leq 2^{O(k)}$ and therefore Theorem 3 is proved.
4.2 Max Ones
Recall that in max ones problem, we given a 3CNF formula $\Phi$ over a variable set $X$ and a weight function $w:X\to\mathbb{R}_{\geq 0}$, the goal is to find a satisfying assignment $\sigma:X\to\{0,1\}$ that maximizes the weight $\sum_{x\in\sigma^{-1}(1)}w(x)$. We reduce the problem to a graph problem. Let $G_{\Phi}=(X\cup\bar{X}\cup\mathcal{C},E)$ be the graph such that $X$ is the variable set of $\Phi$, $\bar{X}=\{\bar{x}_{i}:x_{i}\in X\}$, $\mathcal{C}$ is the clause set of $\Phi$, $\{x_{i},\bar{x}_{i}\}\in E$ for all $x_{i}\in X$, $\{C_{j},x_{i}\}\in E$ iff clause $C_{j}$ contains $x_{i}$ as a positive literal, and $\{C_{j},\bar{x}_{i}\}\in E$ iff clause $C_{j}$ contains $x_{i}$ as a negative literal. Let $w^{\prime}(x)=w(x)$ if $x\in X$, and otherwise $w^{\prime}(x)=0$. Note that the treewidth of $G_{\Phi}$ is at most $2|X|$.
Here, the notations $a\in A\cup B$, $a\in A\cap B$, $a\in A\setminus B$, $\forall a\in A\;\psi$, and $\exists a\in A\;\psi$ are syntactic sugar defined in the usual sense. For a $k$-graph $G$ and $S,D\subseteq\mathsf{src}(G)$, let
$\varphi_{S,D}(A)\equiv\exists B\begin{bmatrix}[l]&\forall a\in A\cup S\;\forall b\in B\;(a\in X\land b\in\bar{X}\land\lnot\mathsf{adj}(a,b))\\ \land&\forall x\in X\cup\bar{X}\;\forall y\in X\cup\bar{X}\;(\mathsf{adj}(x,y)\to(x\in A\cup S\cup B\lor y\in A\cup S\cup B)\\ \land&\forall s\in\mathsf{src}(G)\;(\lnot s\in A)\land\forall c\in\mathcal{C}\setminus D\;\exists x\in A\cup S\cup B\;(\mathsf{adj}(x,c))\end{bmatrix}.$
The first and the second rows says that $A\cup S$ and $B$ represent the sets $\sigma^{-1}(1)$ and $\sigma^{-1}(0)$ of an assignment $\sigma$, respectively. The third row says that $A$ has no sources, and all clauses except in $D$ are satisfied by the $\sigma$ defined by $A\cup S$ and $B$. Then, the max ones problem is equivalent to find a vertex set $Y$ maximizes the weight $\sum_{x\in Y}w^{\prime}(x)$ and satisfies $G_{\Phi}\models\varphi_{\emptyset,\emptyset}(Y)$. Here, we have
$\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi_{S,D})$ =$=\begin{cases}\biguplus_{S^{\prime}\subseteq B}\mathsf{sat}(G,\varphi_{S\cup S^{\prime},D})\boxtimes\{S^{\prime}\}&\text{if $D\cap B=\emptyset$},\\ \emptyset&\text{otherwise}\end{cases}$
$\mathsf{sat}(G\mathbin{//}H,\varphi_{S,D})$ $=\biguplus_{\begin{subarray}{c}D_{1}\cap D_{2}=D\\ D_{1},D_{2}\subseteq\mathsf{src}(G)\end{subarray}}\mathsf{sat}(G,\varphi_{S,D_{1}})\boxtimes\mathsf{sat}(H,\varphi_{S,D_{2}}).$
Particularly, we have $p_{\max}\leq 2^{O(k)}$.
5 Baker’s Technique
In this section, we provide a technique for Lipschitz continuous algorithms for optimization problems on planar graphs. Our technique is a modification of Baker’s technique [1], which provides a PTAS for many optimization problems on planar graphs. For simplicity, in this section, we focus on the maximum weight independent set problem on planar graphs. Similar techniques can be applied to obtain Lipschitz continuous algorithms for other problems, such as the minimum weight vertex cover problem on planar graphs.
Let $G=(V,E)$ be a planar graph and $w\in\mathbb{R}_{\geq 0}^{V}$ be a weight vector. A vertex subset $X$ is an independent set of $G$ if no two vertices in $X$ are adjacent. The maximum weight planar independent set problem asks to maximize the weight $w(X):=\sum_{v\in X}w_{v}$ among all independent sets of $G$. We denote this maximum value by $\mathsf{opt}$. Let $\varepsilon>0$. We provide a $(1-\varepsilon)$-approximation algorithm for the maximum weight planar independent set problem with Lipschitz constant $O((\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n)$.
Let $r\in V$ be an arbitrary vertex. For $i\in\mathbb{Z}_{\geq 0}$, let $R_{i}$ be the set of vertices at exact distance $i$ from $r$, where the distance between $r$ and $v$ is measured as the minimum number of edges in an $r$-$v$ path. Let $m\in\mathbb{Z}_{\geq 1}$. For $j\in\{0,\dots,m-1\}$, let $G_{j}$ be the subgraph of $G$ induced by $V\setminus\bigcup_{i\in\mathbb{Z}_{\geq 0}\colon i\equiv j\pmod{m}}R_{i}$. We denote the maximum weight of an independent set of $G_{j}$ by $\mathsf{opt}_{j}$. The following lemmas are known.
Lemma 25 ([23]).
For each $k\in\{0,\dots,m-1\}$, the treewidth of $G_{i}$ is at most $23m-$ 2.
Lemma 26 ([1]).
We have $\max_{j\in\{0,\dots,m-1\}}\mathsf{opt}_{j}\geq\left(1-\frac{1}{m}\right)\mathsf{opt}$.
In the original Baker’s technique, a $(1-\varepsilon)$-approximation algorithm for the maximum weight planar independent set problem is obtained by finding the maximum weight independent set for each $G_{j}$ and then outputting the one with the maximum weight among them. We make this algorithm Lipschitz continuous by replacing the max operation with a softmax operation. We set $m=\lceil 2\varepsilon^{-1}\rceil$. Then, we take $\bm{j}^{*}=\mathsf{softargmax}^{m^{-1}}_{j\in\{0,\dots,m-1\}}\mathsf{opt}_{j}$ and compute an approximate maximum weight independent set of $G_{\bm{j}^{*}}$ using the algorithm in Section 3, where $\varepsilon$ is set as $m^{-1}$.
Lemma 27.
Our algorithm has approximation ratio $1-\varepsilon$.
Proof.
From Lemma 16, we have $\mathbb{E}\left[\mathsf{opt}_{\bm{j}^{*}}\right]\geq\left(1-\frac{1}{m}\right)\max_{j\in\{0,\dots,m-1\}}\mathsf{opt}_{j}$. For each $j\in\{0,\dots,m-1\}$, our framework outputs $\left(1-\frac{1}{m}\right)$-approximate solution for the maximum weight independent set problem on $G_{j}$. Therefore, the overall approximation ratio is $\left(1-\frac{1}{m}\right)^{2}\geq 1-\frac{2}{m}\geq 1-\varepsilon$. ∎
Lemma 28.
Our algorithm has a Lipschitz constant $O((\varepsilon^{-1}+\log\log n)\varepsilon^{-1}\log^{2}n)$.
Proof.
Let $w\in\mathbb{R}_{\geq 0}^{V}$, $u\in V$, and $\delta>0$. We bound $\mathsf{EM}\left(\bm{X}_{w},\bm{X}_{w+\delta\bm{1}_{u}}\right)$, where $\bm{X}_{w}$ and $\bm{X}_{w+\delta\bm{1}_{u}}$ are the output of our algorithm on weight vector $w$ and $w+\delta\bm{1}_{u}$, respectively. We denote the variable $\bm{j}^{*}$ corresponding to the weight vector $w$ and $w+\delta\bm{1}_{u}$ by $\bm{j}^{*}$ and $\bm{j}^{*}_{w+\delta\bm{1}_{u}}$, respectively. From Lemma 17, we have
$\mathsf{TV}\left(\bm{j}^{*}_{w},\bm{j}^{*}_{w+\delta\bm{1}_{u}}\right)\leq\frac{10m\log(2m\varepsilon^{-1})\delta}{\max_{j\in\{0,\dots,m-1\}}\mathsf{opt}_{j}}\leq\frac{10m\log(2p\varepsilon^{-1})\delta}{\left(1-\frac{1}{m}\right)\mathsf{opt}}.$
Therefore, we have
$\mathsf{EM}\left(\bm{X}_{w},\bm{X}_{w+\delta\bm{1}_{u}}\right)$ $\leq\mathsf{TV}\left(\bm{j}^{*}_{w},\bm{j}^{*}_{w+\delta\bm{1}_{u}}\right)\mathsf{opt}+O\left(m+\log m+\log\log n\right)m\log^{2}n\delta$
$\leq\left(O(m\log m)+O\left(\left(m+\log\log n\right)m\log^{2}n\right)\right)\delta$
$\leq O\left(\left(m+\log\log n\right)m\log^{2}n\right)\delta$
$=O\left(\left(\varepsilon^{-1}+\log\log n\right)\varepsilon^{-1}\log^{2}n\right)\delta,\qed$
where the first inequality is from Theorem 3.
6 Lower Bounds for Max Ones
In this section, we prove Theorem 5. To this end, we use a known sensitivity lower bound for the maximum cut problem. In the maximum cut problem, given a graph $G=(V,E)$, the goal is to find a bipartition of $V$ that maximizes the number of edges between the two parts. Then, the sensitivity of a randomized algorithm $\mathcal{A}$ for the maximum cut problem on a graph $G=(V,E)$ is defined to be
$\max_{e\in E}\mathsf{EM}(\mathcal{A}(G),\mathcal{A}(G-e)),$
where $G-e$ denotes the graph obtained from $G$ by deleting $e$ and the distance $d$ underlying the earth mover’s distance is the Hamming distance, i.e., $d(S,S^{\prime})=|S\triangle S^{\prime}|$. We use the following lower bound:
Lemma 29 ([9]).
There exist $\varepsilon,\delta>0$ such that any $(1-\varepsilon)$-approximation algorithm for the maximum cut problem sensitivity $\Omega(n^{\delta})$, even on constant-degree graphs.
Proof of Theorem 5.
Let $\varepsilon,\delta>0$ be as in Lemma 29. Assume that there exists a $(1-\varepsilon)$-approximation algorithm $\mathcal{A}_{\rm MO}$ for the max ones problem with Lipschitz constant $s(n)$, where $n$ is the number of variables in the input 3CNF formula. Below, we design a low sensitivity algorithm $\mathcal{A}_{\rm MC}$ for the maximum cut problem using $\mathcal{A}_{\rm MO}$.
Given an unweighted graph $G=(V,E)$, we construct a 3CNF formula $\varphi$ over $X:=\{x_{v}:v\in V\}\cup\{x_{uv}:\{u,v\}\in\binom{V}{2}\}$ as follows. For each pair $\{u,v\}\in\binom{V}{2}$, we add constraints over $x_{u}$, $x_{v}$, and $x_{uv}$ that are satisfied if and only if (i) $x_{uv}=1$ and $x_{u}\neq x_{v}$ or (ii) $x_{uv}=0$. More specifically, we add constraints $(\bar{x}_{uv}\vee x_{u}\vee x_{v})$ and $(\bar{x}_{uv}\vee\bar{x}_{u}\vee\bar{x}_{v})$. Note that $\varphi$ does not depend on $E$. Next, we define a weight function $w:X\to\mathbb{R}_{\geq 0}$ such that $w(x_{u})=0$ for every $u\in V$ and $w(x_{uv})=1$ if and only if $\{u,v\}\in E$. Then, we run the algorithm $\mathcal{A}_{\rm MO}$ on the instance $(\varphi,w)$, and let $\sigma:X\to\{0,1\}$ be the obtained solution. Finally, let $S:=\sigma^{-1}(1)\cap V$ be the solution for $G$, and outputs $S$ with probability half and $V\setminus S$ with the remaining probability of half.
Note that the total weight of edges cut by $S$ in $G$ is equal to the total weight of $\sigma^{-1}(1)$. Hence, $\mathcal{A}_{\rm MC}$ is a $(1-\varepsilon)$-approximation algorithm for the maximum cut problem.
Now, we analyze the sensitivity of $\mathcal{A}_{\rm MC}$. Let $G=(V,E)$ and $G^{\prime}=(V,E^{\prime})$ be two graphs with $|E\triangle E^{\prime}|=1$. Let $w$ and $w^{\prime}$ be the weight functions constructed from $G$ and $G^{\prime}$, respectively. Note that $\|w-w^{\prime}\|_{1}=1$. Let $\sigma$ and $\sigma^{\prime}$ be the outputs of $\mathcal{A}_{\rm MO}$ on $(\varphi,w)$ and $(\varphi,w^{\prime})$, respectively. Let $F=\{\{u,v\}\in E\mid\sigma(x_{uv})=1\}$ and $F^{\prime}=\{\{u,v\}\in E^{\prime}\mid\sigma^{\prime}(x_{uv})=1\}$. Let $S$ and $S^{\prime}$ be the outputs of $\mathcal{A}_{\mathrm{M}C}$ for $G$ and $G^{\prime}$, respectively. Then, we have
$\mathsf{EM}(S,S^{\prime})\leq\Delta\cdot\mathsf{EM}(F,F^{\prime})=O(\mathsf{EM}(F,F^{\prime})),$
where $\Delta=O(1)$ is the degree bound of the graphs $G$ and $G^{\prime}$. Because $\mathcal{A}_{\rm MO}$ has Lipschitz constant $s(n)$ and the number of variables in $\varphi$ is $n^{3}$, we have
$\mathsf{EM}(S,S^{\prime})\leq O(\mathsf{EM}(F,F^{\prime}))=O(\mathsf{EM}((\sigma^{-1}(1),w),((\sigma^{\prime})^{-1}(1),w^{\prime})))$
$=O\left(\frac{\mathsf{EM}((\sigma,w),(\sigma^{\prime},w^{\prime}))}{\|w-w^{\prime}\|_{1}}\right)=O(s(n^{3})).$
This implies that $\mathcal{A}_{\rm MC}$ has sensitivity at most $O(s(n^{3}))$. By Lemma 29, we must have $s(n)=\Omega(n^{\delta/3})$. ∎
7 Clique-Width
In this section, we prove Theorems 6 and 7. Since the proofs are similar to that for Theorem 6, we provide most of the proof of Theorem 7 in Appendix C. Let $G$ be a vertex-colored graph with $n$ vertices. A VR-term $\mathcal{T}$ of $G$ represents a procedure that constructs $G$ by recursively applying the following operations [4]:
  • Create a vertex with color $i$.
  • Take the disjoint union of two vertex-colored graphs with VR-terms $T_{1},T_{2}$.
  • Choose two different colors $i,j$ and add edges between all vertices with color $i$ and all vertices with color $j$.
  • Choose two different colors $i,j$ and recolor all vertices with color $i$ by color $j$.
The width of a VR-term is the number of distinct colors that appear in the overall procedure. For a vertex $v$ in $G$, the height of $v$ with respect to a VR-term $\mathcal{T}$ is the number of times that the disjoint union operation is applied to a graph containing $v$ during the construction procedure. The height of $\mathcal{T}$ is the maximum height among all vertices with respect to $\mathcal{T}$. The clique-width of a graph $G$ is the minimum width among all VR-terms of $G$. We have the following.
Lemma 30 ([5, 21]).
Let $G$ be a graph with $n$ vertices and $\mathrm{cw}$ be the clique-width of $G$. Then, a VR-term of $G$ with width $O^{*}(2^{\mathrm{cw}})$ and height $O(\log n)$ can be computed in $2^{O(\mathrm{cw})}n^{3}$ time.
Theorem 31 ([4]).
Let $r,q,k$ be positive integers. Let $\varphi({X})$ be a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula over $k$-colored graph with a second-order free variable ${X}$. Then, the following holds.
  1. 1.
    For any $i,j\in[k]$, there exist $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas $\psi({X})$ and $\psi^{\prime}({X})$ such that for any colored graph $G$, we have
    $\mathsf{sat}(\eta_{i,j}(G),\varphi)=\mathsf{sat}(G,\psi),\quad\mathsf{sat}(\rho_{i\to j}(G),\varphi)=\mathsf{sat}(G,\psi^{\prime}),$
    where, $\eta_{i,j}(G)$ is the graph obtained from $G$ by adding edges between all vertices with color $i$ and all vertices with color $j$, $\rho_{i\to j}(G)$ is the graph that obtained from $G$ by recoloring all vertices with color $i$ by color $j$.
  2. 2.
    There exists a family of tuples $\{\theta_{i}({X}),\psi_{i}({X})\}_{i\in[p]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with a free variable ${X}$ such that, for any $k$-colored $G$ and $H$,
    $\mathsf{sat}(G\oplus H,\varphi)=\biguplus_{i\in[p]}\mathsf{sat}(G,\theta_{i})\boxtimes\mathsf{sat}(H,\psi_{i}).$
7.1 Definitions
For a graph $G$, a weight vector $w$, and an $\textsf{MSO}_{1}$ formula $\varphi$, we define
$\mathsf{opt}_{w}[G,\varphi]=\max_{G\models\varphi(S)}w(S).$ (5)
for maximization problems and
$\mathsf{opt}_{w}[G,\varphi]=\min_{G\models\varphi(S)}w(S).$ (6)
for minimization problems. If $\mathsf{sat}(G,\varphi)=\emptyset$, we define $\mathsf{opt}_{w}[G,\varphi]=\mathtt{nil}$. For graphs $G$ and $\textsf{MSO}_{1}$ formulas $\varphi$ with $\mathsf{sat}(G,\varphi)\neq\emptyset$, our algorithm recursively computes the vertex set $\mathsf{DP}_{w}[G,\varphi]$ that approximately achieves the maximum in Equation (5) or minimum in Equation (6). When it is clear from the context, we omit the subscript $w$ and simply write $\mathsf{opt}[G,\varphi]$ and $\mathsf{DP}[G,\varphi]$. We perform the dynamic programming algorithm using the formulas defined in Theorem 31 over the parse tree of a term obtained from Lemma 30. Specifically, at every stage of the algorithm, it is guaranteed that each $X\in\mathsf{sat}(G,\varphi)$ satisfies $X\subseteq V(G)$. Furthermore, our algorithm is randomized. Thus, $\mathsf{DP}[G,\varphi]$ can be considered as a probability distribution over vertex sets in $\mathsf{sat}(G,\varphi)$.
To bound the approximation ratio, for a weight vector $w\in\mathbb{R}_{\geq 0}^{V}$, we bound the ratio between $\mathbb{E}\left[w(\mathsf{DP}[G,\varphi])\right]$ and $\mathsf{opt}[G,\varphi]$. To bound the Lipschitz constant, for a weight vector $w\in\mathbb{R}_{\geq 0}^{V}$, $u\in V$, and $\delta>0$, we bound $\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi]\right)$. From now on, we will concentrate on maximizing problems. The algorithm for minimization problems is similar to that for maximization problems, while the detail of the analysis is slightly different. We discuss the minimization version in Appendix C.
7.2 Base Case
[…]
7.3 Disjoint Union
Next, we consider the disjoint union. The algorithm and discussion here are the same as those in Section 3.3, where the parallel composition is replaced by the disjoint union. Therefore, we omit the technical details and only provide the definition of the DP table, as well as the bounds on approximability and Lipschitz continuity. Let
$\mathsf{sat}(G\oplus H,\varphi)=\biguplus_{i\in[p]}\mathsf{sat}(G,\theta_{i})\boxtimes\mathsf{sat}(H,\psi_{i}).$ (7)
Assume we have already computed $\mathsf{DP}[G,\theta_{i}]$ and $\mathsf{DP}[H,\psi_{i}]$ for each $i\in[p]$. For each $i\in[p]$, we denote $\mathsf{opt}_{i}:=\mathsf{opt}[G,\theta_{i}]+\mathsf{opt}[H,\psi_{i}]$. Then, we take $\bm{i}^{*}=\mathsf{softargmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}$ and define $\mathsf{DP}[G\oplus H,\varphi]:=\mathsf{DP}[G,\theta_{\bm{i}^{*}}]\cup\mathsf{DP}[H,\psi_{\bm{i}^{*}}]$. We have the following lemmas.
Lemma 32.
Let $0<\alpha\leq 1$ and suppose $\mathbb{E}\left[w(\mathsf{DP}[G,\theta_{i}])\right]\geq\alpha\mathsf{opt}[G,\theta_{i}]$ and $\mathbb{E}\left[w(\mathsf{DP}[H,\psi_{i}])\right]\geq\alpha\mathsf{opt}[H,\psi_{i}]$ hold for all $i\in[p]$. Then, we have $\mathbb{E}\left[w\left(\mathsf{DP}[G\oplus H,\varphi]\right)\right]\geq(1-\varepsilon)\alpha\mathsf{opt}[G\oplus H,\varphi]$.
Lemma 33.
Let $\beta\in\mathbb{R}_{\geq 0}$ and suppose $\mathsf{EM}\left(\mathsf{DP}_{w}[G,\theta_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\theta_{i}]\right)\leq\beta$ holds for all $i\in[p]$. Then, we have $\mathsf{EM}\left(\mathsf{DP}_{w}[G\oplus H,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G\oplus H,\varphi]\right)\leq 30\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta$.
7.4 Putting Together
Here we complete the proof of Theorem 6. Let $G=(V,E)$ be a graph with clique-width $\mathrm{cw}$. From Lemma 8, we can compute a VR-term $\mathcal{T}$ over a $k$-colored graph denoting $G$ with height $O(\log n)$. We perform the dynamic programming algorithm on the parse tree of $\mathcal{T}$. We remark here that operations other than disjoint union do not affect either the approximability or the Lipschitz continuity, by setting
$\mathsf{DP}[\eta_{i,j}(G),\varphi]:=\mathsf{DP}[G,\psi],\quad\mathsf{DP}[\rho_{i\to j}(G),\varphi]:=\mathsf{DP}[G,\psi^{\prime}].$
We have the following.
Lemma 34.
Let $\varepsilon\in(0,1]$ and $h$ be the height of the VR-term $\mathcal{T}$. Our algorithm outputs a solution $X\in\mathsf{sat}(G,\varphi)$ such that $\mathbb{E}[w(X)]\geq(1-h\varepsilon)\mathsf{opt}[G,\varphi]$. The Lipschitz constant is bounded by $30(h+1)\varepsilon^{-1}\log(2p_{\max}\varepsilon^{-1})$, where $p_{\max}$ is the maximum of $p$ among all update formula (7) that the algorithm uses.
Proof.
The approximability bound is obtained by repeatedly applying Lemma 32 and $(1-\varepsilon)^{h}\geq 1-h\varepsilon$. The Lipschitzness bound is obtained by repeatedly applying Lemma 33. ∎
Since $p_{\max}$ is bounded by a function of $k$ and $|\varphi|$, we have the following.
Proof of Theorem 6.
The claim follows by substituting $\varepsilon$, $h$, and $k$ in Lemma 24 with $\frac{\varepsilon}{\Theta(\log n)}$, $O(\log n)$, and $O^{*}(2^{k})$, respectively. ∎
Appendix A Omitted Proofs on Logic
Proof of Proposition 11.
Let $(\mathcal{B},\mathcal{T})$ be a tree decomposition of $G$. Let $t\in V(\mathcal{T})$. Let $\mathcal{T}_{t}$ denote the subtree of $\mathcal{T}$ consisting of all descendants of $t$. Note that $\mathcal{T}_{t}\subseteq V(G)$ since a node of $\mathcal{T}$ is in $\mathcal{B}$.
We recursively transform a tree decomposition $(\mathcal{B},\mathcal{T})$ to a HR-parse tree. In this proof, we denote the height of the parse tree of a term $t$ by $0pt(t)$. Let $v=B_{v}$ be a node of $\mathcal{T}$. We have the following three cases.
$v$ is a leaf node:
For a graph $H$, let $\mathtt{Graph}(H)$ denote the term $(\mathbin{//}_{i\in V(H)}\mathtt{i})\mathbin{//}(\mathbin{//}_{\{i,j\}\in E(H)}\mathtt{ij})$. We arbitrarily identify $B_{v}$ with $\{1,\ldots,|B_{v}|\}$. Then, it is clear that the term $\mathtt{Graph}(G[B_{v}])$ denote the graph $G[B_{v}]$. When $\mathtt{Graph}(G[B_{v}])$ is expanded linearly, its height is $O(|B_{v}|^{2})=O(k^{2})$. However, by restructuring it to a balanced binary tree, the height can be reduced to $O(\log k)$ since parallel-composition operation has the associative property.
$v$ has exactly one child node $c=B_{c}$:
Let $t_{c}$ be the term representing $G[\mathcal{T}_{c}]$. Let $G_{\Delta}=G[B_{v}\setminus B_{c}]$. Then, the term $\mathtt{Graph}(G_{\Delta})\mathbin{//}\mathsf{fg}_{B_{c}\setminus B_{v}}(t_{c})$ denotes the graph $G[\mathcal{T}_{v}]$ and its height is at most $\max\{\log(k),0pt(t_{c})\}+2$.
$v$ has exactly two child nodes $c_{1}=B_{c_{1}}$ and $c_{2}=B_{c_{2}}$:
Let $t_{1}$ and $t_{2}$ be the terms representing $G[\mathcal{T}_{c_{1}}]$ and $G[\mathcal{T}_{c_{2}}]$, respectively. Let $G_{\Delta}=G[B_{v}\setminus(B_{c_{1}}\cup B_{c_{2}})]$. Then, the term $\mathtt{Graph}(G_{\Delta})\mathbin{//}\mathsf{fg}_{B_{c}\setminus B_{v}}(t_{1})\mathbin{//}\mathsf{fg}_{B_{c_{2}}\setminus B_{v}}(t_{2})$ denotes the graph $G[\mathcal{T}_{v}]$ and the height is at most $\max\{\log(k),\allowbreak 0pt(t_{1}),0pt(t_{2})\}+3$.
It is easy to verify that the resulting parse tree represents $G$, and the height is at most $O(\log k+h)$. ∎
Proof of Corollary 14.
First we show the first statement of Corollary 14. Since $S^{\prime}\subseteq B\subseteq\mathsf{src}(G)$, $\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})$ and $\{S^{\prime}\}$ is separated. It is clear that families of vertex sets $(\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\})_{S^{\prime}\subseteq B}$ are disjoint.
Let $\psi(X)$ be a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula obtained by applying the first statement of Theorem 12 to $\varphi(X)$ and $B$. Then, $\mathsf{fg}_{B}(G)\models\varphi(A)\iff G\models\psi(A)$ for any vertex set $A$. We show that $\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi\mathord{{\restriction}_{S}})=\biguplus_{S^{\prime}\subseteq B}\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\}.$
Let $A\in\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi\mathord{{\restriction}_{S}})$ and $A^{\prime}=(\mathsf{src}(G)\cap A)\setminus S$. Since $\mathsf{fg}_{B}(G)\models\varphi\mathord{{\restriction}_{S}}(X)$, we have $\mathsf{fg}_{B}(G)\models\varphi(A\cup S)$ and $\mathsf{fg}_{B}(G)\models\mathsf{nosrc}(A)$. Thus, $G\models\psi(A\cup S)$ and then $G\models\psi((A\setminus A^{\prime})\cup(A\cup A^{\prime}))$. Furthermore, $A\cap\mathsf{src}({\mathsf{fg}_{B}(G)})=\varnothing$ and thus,
$A^{\prime}=(\mathsf{src}(G)\cap A)\setminus S=big((\mathsf{src}(G)\setminus\mathsf{src}(\mathsf{fg}_{B}(G)))\cap Abig)\setminus S=(B\cap A)\setminus S.$
Since $B\cap S=\varnothing$, we have $A^{\prime}=B\cap A$, and thus, $G\models\mathsf{nosrc}(A\setminus A^{\prime})$. Thus, $A\in\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup A^{\prime}}})\boxtimes\{A^{\prime}\}$ and $A^{\prime}\subseteq B$.
Let $A\in\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\}$ for some $S^{\prime}\subseteq B$. Then, $G\models\psi((A\setminus S^{\prime})\cup S\cup S^{\prime})$. Thus, $\mathsf{fg}_{B}(G)\models\varphi(A\cup S)$. Since $G\models\mathsf{nosrc}(A\setminus S^{\prime})$ and $\mathsf{src}(\mathsf{fg}_{B}(G))=\mathsf{src}(G)\setminus B\subseteq\mathsf{src}(G)\setminus S^{\prime}$, we have $A\cap\mathsf{src}(\mathsf{fg}_{B}(G))=\varnothing$. Thus, $A\in\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi\mathord{{\restriction}_{S}})$.
Next, we show the second statement of Corollary 14. From Theorem 12, there exists a family of tuples $\{\theta^{\prime}_{i}({X}),\psi^{\prime}_{i}({X})\}_{i\in[p^{\prime}]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with a free variable ${X}$ such that, for any $k$-graphs $G$ and $H$,
$\mathsf{sat}(G\mathbin{//}H,\varphi)=\biguplus_{i\in[p^{\prime}]}\{S\cup P:S\in\mathsf{sat}(G,\theta^{\prime}_{i}),P\in\mathsf{sat}(H,\psi^{\prime}_{i})\}.$
Then, we have
$\mathsf{sat}$ $(G\mathbin{//}H,\varphi\mathord{{\restriction}_{S}})\boxtimes\{S\}$
$=\mathsf{sat}(G\mathbin{//}H,\varphi)\cap\{X\colon\mathsf{src}(G\mathbin{//}H)\cap X=S\}$
$=\Bigg{(}\biguplus_{i\in[p]}\left\{S\cup P:S\in\mathsf{sat}(G,\theta^{\prime}_{i}),P\in\mathsf{sat}(H,\psi^{\prime}_{i})\right\}\Bigg{)}\cap\{X\colon\mathsf{src}(G\mathbin{//}H)\cap X=S\}$
[…]
$=\biguplus_{i\in[p]}\bigcup_{\begin{subarray}{c}(S^{\prime},S^{\prime\prime})\colon\\ S^{\prime}\cup S^{\prime\prime}=S,\\ S^{\prime}\subseteq\mathsf{src}(G),\ S^{\prime\prime}\subseteq\mathsf{src}(H)\end{subarray}}\mathsf{sat}(G,\theta_{i}\mathord{{\restriction}_{S^{\prime}}})\boxtimes\mathsf{sat}(H,\psi_{i}\mathord{{\restriction}_{S^{\prime\prime}}})\boxtimes\{S\}.$
Thus, there exists a family of tuples $\{\theta_{i}\mathord{{\restriction}_{S_{i}}}({X}),\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}({X})\}_{i\in[p]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with free variables ${X}$ such that, for any $k$-graphs $G$ and $H$,
$\mathsf{sat}(G\mathbin{//}H,\varphi\mathord{{\restriction}_{S}})=\bigcup_{i\in[p^{\prime}]}\mathsf{sat}(G,\theta_{i}\mathord{{\restriction}_{S_{i}}})\boxtimes\mathsf{sat}(H,\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}).$
[…]
Appendix B Analysis of Softmax and Softmin
In this section, we provide the approximation ratios and stabilities of $\mathsf{softmax}$ and $\mathsf{softmin}$. The following lemma is useful.
Lemma 35.
Let $x\geq\delta>0$ and $q>0$. Consider sampling $\bm{c}$ uniformly from $\left[\frac{q}{x},\frac{2q}{x}\right]$ and $\bm{c}^{\prime}$ uniformly from $\left[\frac{q}{x+\delta},\frac{2q}{x+\delta}\right]$. Then, we have $\mathsf{TV}(\bm{c},\bm{c}^{\prime})\leq\frac{2\delta}{x}$.
Proof.
We have
$\mathsf{TV}\left(\bm{c},\bm{c}^{\prime}\right)=\Pr\left[\bm{c}\in\left[\frac{2q}{x+\delta},\frac{2q}{x}\right]\right]=\frac{\frac{2q}{x}-\frac{2q}{x+\delta}}{\frac{2q}{x}-\frac{q}{x}}=2-\frac{2x}{x+\delta}=\frac{2\delta}{x+\delta}\leq\frac{2\delta}{x}.\qed$
Now we analyze the $\mathsf{softmax}$ operation.
Proof of Lemma 16.
We have
$\mathbb{E}\left[\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}\right]$ $\geq\Pr\left[\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}\geq\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\right]\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}$
$\geq\left(1-p\cdot\frac{\exp\left(c\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\right)}{\sum_{i\in[p]}\exp(cx_{i})}\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}$
$\geq\left(1-p\cdot\frac{\exp\left(c\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\right)}{\max_{i\in[p]}\exp(cx_{i})}\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}$
$=\left(1-p\cdot\frac{\exp\left(c\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\right)}{\exp\left(c\max_{i\in[p]}x_{i}\right)}\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}$
$=\left(1-p\cdot\exp\left(-\frac{c\epsilon}{2}\max_{i\in[p]}x_{i}\right)\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}$
$\geq\left(1-p\cdot\exp\left(\log(2p\epsilon^{-1})\right)\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}$
$=\left(1-\frac{\epsilon}{2}\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\geq(1-\epsilon)\max_{i\in[p]}x_{i}.\qed$
Proof of Lemma 17.
Let $\bm{c}$ and $\bm{c}^{\prime}$ be the vectors sampled when computing $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}$ and $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}$, respectively. Let $I=\left[\frac{2\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}},\frac{4\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}^{\prime}}\right]$. For each parameter $c\in I$, we consider transporting probability mass for $\bm{c}=c$ to that for $\bm{c}^{\prime}=c$ as far as possible. The remaining mass is transported arbitrarily. Then, we have
$\mathsf{TV}\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)$
$\leq\mathsf{TV}\left(\bm{c},\bm{c}^{\prime}\right)+\sup_{c\in I}\mathsf{TV}\left(\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}\mid\bm{c}=c\right),\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\mid\bm{c^{\prime}}=c\right)\right).$
Now, for $c\in I$, we have
$\mathsf{TV}\left(\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}\mid\bm{c}=c\right),\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\mid\bm{c^{\prime}}=c\right)\right)$
$=\sum_{i\in[p]}\max\left(0,\Pr\left[\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}=i\mid\bm{c}^{\prime}=c\right]-\Pr\left[\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}=i\mid\bm{c}=c\right]\right)$
$=\sum_{i\in[p]}\max\left(0,\frac{\exp(cx_{i}^{\prime})}{\sum_{i^{\prime}\in[p]}\exp(cx_{i^{\prime}}^{\prime})}-\frac{\exp(cx_{i})}{\sum_{i^{\prime}\in[p]}\exp(cx_{i^{\prime}})}\right)$
$\leq\sum_{i\in[p]}\frac{\exp(cx_{i}^{\prime})-\exp(cx_{i})}{\sum_{i^{\prime}\in[p]}\exp(cx_{i^{\prime}})}=\sum_{i\in[p]}\frac{(\exp(c(x_{i}^{\prime}-x_{i}))-1)\exp(cx_{i})}{\sum_{i^{\prime}}\exp(cx_{i^{\prime}})}$
$\leq\sum_{i\in[p]}\frac{(\exp(c\delta)-1)\exp(cx_{i})}{\sum_{i^{\prime}}\exp(cx_{i^{\prime}})}=\exp(c\delta)-1\leq 2c\delta\leq\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}},$
where the fourth inequality is from the fact that $\exp(x)-1\leq 2x$ holds for $x\leq 1$. Therefore, we have
$\mathsf{TV}\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{2\delta}{\max_{i\in[p]}x_{i}}+\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}.\qed$
Now we analyze the $\mathsf{softmin}$ operation.
Proof of Lemma 18.
For $t\geq\frac{1}{2}$, we have
$\Pr\left[x_{\bm{i}^{*}}\geq\left(1+t\epsilon\right)\min_{i\in[p]}x_{i}\right]$ $\leq\frac{p\exp(-c(1+t\epsilon)\min_{i\in[p]}x_{i})}{\sum_{i\in[p]}\exp(-cx_{i})}\leq\frac{p\exp(-c(1+t\epsilon)\min_{i\in[p]}x_{i})}{\max_{i\in[p]}\exp(-cx_{i})}$
$=\frac{p\exp(-c(1+t\epsilon)\min_{i\in[p]}x_{i})}{\exp(-c\min_{i\in[p]}x_{i})}=p\exp\left(-ct\epsilon\min_{i\in[p]}x_{i}\right)$
$=p\exp(-tb\log(2p\epsilon^{-1}))\leq p\cdot\left(\frac{\epsilon}{2p}\right)^{2t}\leq\left(\frac{\epsilon}{2}\right)^{2t}.$
Therefore, we have
[…]
$\leq\left(\left(1+\frac{\epsilon}{2}\right)+\left(\frac{\epsilon}{2}\right)^{2+\epsilon}\cdot 2\log\left(2\epsilon^{-1}\right)\right)\cdot\min_{i\in[p]}x_{i}$
$\leq\left(\left(1+\frac{\epsilon}{2}\right)+\left(\frac{\epsilon}{2}\right)^{2}\cdot 2\right)\cdot\min_{i\in[p]}x_{i}\leq(1+\epsilon)\min_{i\in[p]}x_{i}.\qed$
Proof of Lemma 19.
Let $\bm{c}$ and $\bm{c}^{\prime}$ be the vectors sampled when computing $\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}$ and $\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}$, respectively. Let $I=\left[\frac{2\epsilon^{-1}\log(2p\epsilon^{-1})}{\min_{i\in[p]}x_{i}},\frac{4\epsilon^{-1}\log(2p\epsilon^{-1})}{\min_{i\in[p]}x_{i}^{\prime}}\right]$. For each parameter $c\in I$, we consider transporting probability mass for $\bm{c}=c$ to that for $\bm{c}^{\prime}=c$ as far as possible. The remaining mass is transported arbitrarily. Then, we have
$\mathsf{TV}\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)$
$\leq\mathsf{TV}\left(\bm{c},\bm{c}^{\prime}\right)+\sup_{c\in I}\mathsf{TV}\left(\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}\mid\bm{c}=c\right),\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\mid\bm{c^{\prime}}=c\right)\right).$
Now, for $c\in I$, we have
$\mathsf{TV}\left(\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}\mid\bm{c}=c\right),\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\mid\bm{c^{\prime}}=c\right)\right)$
$=\sum_{i\in[p]}\max\left(0,\Pr\left[\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}=i\mid\bm{c}=c\right]-\Pr\left[\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}=i\mid\bm{c}^{\prime}=c\right]\right)$
$=\sum_{i\in[p]}\max\left(0,\frac{\exp(-cx_{i})}{\sum_{i^{\prime}\in[p]}\exp(-cx_{i^{\prime}})}-\frac{\exp(-cx_{i}^{\prime})}{\sum_{i^{\prime}\in[p]}\exp(-cx_{i^{\prime}}^{\prime})}\right)$
$\leq\sum_{i\in[p]}\frac{\exp(-cx_{i})-\exp(-cx_{i}^{\prime})}{\sum_{i^{\prime}\in[p]}\exp(-cx_{i^{\prime}})}=\sum_{i\in[p]}\frac{(1-\exp(-c(x_{i}^{\prime}-x_{i})))\exp(-cx_{i})}{\sum_{i^{\prime}\in[p]}\exp(-cx_{i^{\prime}})}$
$\leq\sum_{i\in[p]}\frac{(1-\exp(-c\delta))\exp(-cx_{i})}{\sum_{i^{\prime}\in[p]}\exp(-cx_{i^{\prime}})}=1-\exp(-c\delta)\leq 2c\delta\leq\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}},$
where the fourth inequality is from the fact that $1-\exp(-x)\leq x\leq 2x$ holds for $x\leq 1$. Therefore, we have
$\mathsf{TV}\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{2\delta}{\min_{i\in[p]}x_{i}}+\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}.\qed$
Appendix C Minimization
C.1 Treewidth
Here, we provide the minimization version of the framework from Section 3. The algorithm is obtained simply by replacing $\mathsf{softargmax}$ with $\mathsf{softargmin}$. Some parts of the analysis are identical to that in Section 3. Specifically, for the introduce operation, exactly the same results can be obtained with exactly the same proof. Furthermore, by replacing $(1+\epsilon)$ with $(1-\epsilon)$ and reversing all inequalities in the proofs of Lemmas 20 and 22, we obtain the following.
Lemma 36.
Let $0<\alpha\leq 1$ and assume
$\mathbb{E}\left[w(\mathsf{DP}[G,\varphi^{G}_{i}])\right]$ $\leq\alpha\mathsf{opt}[G,\varphi^{G}_{i}],$
$\mathbb{E}\left[w(\mathsf{DP}[H,\varphi^{H}_{i}])\right]$ $\leq\alpha\mathsf{opt}[H,\varphi^{H}_{i}]$
holds for all $i\in[p]$. Then, we have $\mathbb{E}\left[w\left(\mathsf{DP}[G\mathbin{//}H,\varphi]\right)\right]\leq(1+\epsilon)\alpha\mathsf{opt}[G\mathbin{//}H,\varphi]$.
Lemma 37.
Let $0<\alpha\leq 1$ and assume
$\mathbb{E}\left[w(\mathsf{DP}[G,\varphi_{i}])\right]$ $\leq\alpha\mathsf{opt}[G,\varphi_{i}]$
holds for all $i\in[p]$. Then, we have $\mathsf{DP}[\mathsf{fg}_{B}(G),\varphi]\leq(1+\epsilon)\alpha\mathsf{opt}[\mathsf{fg}_{B}(S),\varphi]$.
In the analysis of Lipschitz continuity, $\mathsf{opt}$ cannot be used as a trivial bound for the earth mover’s distance. Thus, we need a slightly different proof.
Lemma 38.
Let $0<\alpha\leq 1$, $\beta\in\mathbb{R}_{\geq 0}$ and assume
$\mathbb{E}\left[w(\mathsf{DP}[G\mathbin{//}H,\varphi])\right]$ $\leq\alpha\mathsf{opt}_{w}[G\mathbin{//}H,\varphi],$
$\mathbb{E}\left[(w+\delta\bm{1}_{u})(\mathsf{DP}[G\mathbin{//}H,\varphi])\right]$ $\leq\alpha\mathsf{opt}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi],$
$\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi^{G}_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi^{G}_{i}]\right)$ $\leq\beta$
holds for all $i\in[p]$. Then, we have
$\mathsf{EM}\left(\mathsf{DP}_{w}[G\mathbin{//}H,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)\leq 30\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.$
Proof.
We have
$\mathsf{EM}\left(\mathsf{DP}_{w}[G\mathbin{//}H,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)$
$\leq\mathsf{TV}\left(\bm{i}^{*}_{w},\bm{i}^{*}_{w+\delta\bm{1}_{u}}\right)\alpha\left(\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)+\max_{i\in[p]}\left(\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi^{G}_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi^{G}_{i}]\right)\right)$
$\leq\frac{10\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\mathsf{opt}_{w}[G\mathbin{//}H,\varphi,S]}\left(\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[G\mathbin{//}H,\varphi]\right)+\beta\leq 30\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.$
Lemma 39.
Let $\beta\in\mathbb{R}_{\geq 0}$ and assume
$\mathbb{E}\left[w(\mathsf{DP}[\mathsf{fg}_{B}(G),\varphi])\right]$ $\leq\alpha\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi],$
$\mathbb{E}\left[(w+\delta\bm{1}_{u})(\mathsf{DP}[\mathsf{fg}_{B}(G),\varphi])\right]$ $\leq\alpha\mathsf{opt}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi],$
$\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi_{i}]\right)\leq\beta$
holds for all $i\in[p]$. Then, we have
$\mathsf{EM}\left(\mathsf{DP}_{w}[\mathsf{fg}_{B}(G),\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)\leq 31\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.$
Proof.
We have
$\mathsf{EM}\left(\mathsf{DP}_{w}[\mathsf{fg}_{B}(G),\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)$
$\leq\mathsf{TV}\left(\bm{i}^{*}_{w},\bm{i}^{*}_{w+\delta\bm{1}_{u}}\right)\alpha\left(\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)$
$\quad+\max_{i\in[p]}\left(\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi_{i}]\cup S_{i},\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi_{i}]\cup S_{i}\right)\right)$
$\leq\frac{10\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi]}\left(\mathsf{opt}_{w}[\mathsf{fg}_{B}(G),\varphi]+\mathsf{opt}_{w+\delta\bm{1}_{u}}[\mathsf{fg}_{B}(G),\varphi]\right)$
$\quad+\max_{i\in[p]}\left(\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi_{i}]\right)\right)+\max_{i\in[p]}\left[d_{\mathrm{w}}((S_{i},w),(S_{i},w+\delta\bm{1}_{u}))\right]$
$\leq 30\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta+\delta\leq 31\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.\qed$
The following lemma corresponds to Lemma 24, but slightly different.
Lemma 40.
Let $\epsilon\in\mathbb{R}_{>0}$ and $h$ be the height of $T$. Our algorithm outputs a solution $X\in\mathsf{sat}(G,\varphi)$ such that $\mathbb{E}[w(X)]\leq(1+\epsilon)^{h}\mathsf{opt}[G,\varphi]$. The Lipschitz constant is bounded by $31h(1+\epsilon)^{h}\epsilon^{-1}\log(g(t,|\varphi|)\epsilon^{-1})$ for some computable function $g$.
Proof.
The approximability bound is obtained by repeatedly applying Lemmas 36 and 37. The Lipschitzness bound is obtained by repeatedly applying Lemmas 38 and 39 together with $\alpha\leq(1+\epsilon)^{h}$. ∎
Thus, we have the following.
Proof of Theorem 2.
We prove the theorem by substituting $\epsilon$, $h$, and $t$ in Lemma 40 with $\frac{\epsilon}{O(\log n)}$, $O(\log n)$, and $23t-$ 2, respectively. Let $c\in\mathbb{R}_{>0}$ be the number such that $c\log n$ is an upper bound of $h$. We substitute $\epsilon$ by $\frac{\epsilon}{2c\log n}$. Then, the approximation ratio is bounded as
$\left(1+\frac{\epsilon}{2c\log n}\right)^{c\log n}\leq\lim_{n\to\infty}\left(1+\frac{\epsilon}{2n}\right)^{n}\leq e^{\frac{\epsilon}{2}}\leq 1+\epsilon,$
where the last inequality is from $\epsilon\in(0,1]$. The Lipschitz constant is bounded as
$31c\log n\left(1+\frac{\epsilon}{2c\log n}\right)^{c\log n}\cdot 2c\log n\epsilon^{-1}\log\left(g(3t-2,|\varphi|)\cdot 2c\log n\epsilon^{-1}\right)$
$\leq O\left(\epsilon^{-1}\log^{2}n\log(g(3t-2,|\varphi|)\epsilon^{-1}\log n)\right)\leq O\left((f(t,|\varphi|)+\log\epsilon^{-1}+\log\log n)\epsilon^{-1}\log^{2}n\right),$
where $f$ is some computable function. ∎
C.2 Clique-Width
Here, we provide the minimization version of the framework from Section 7. Like in Section 7, the discussion and proof are the same as in Section C.1. Therefore, we only provide the statements of the following lemmas.
Lemma 41.
Let $0<\alpha\leq 1$ and assume
$\mathbb{E}\left[w(\mathsf{DP}[G,\varphi^{G}_{i}])\right]$ $\leq\alpha\mathsf{opt}[G,\varphi^{G}_{i}],$
$\mathbb{E}\left[w(\mathsf{DP}[H,\varphi^{H}_{i}])\right]$ $\leq\alpha\mathsf{opt}[H,\varphi^{H}_{i}]$
holds for all $i\in[p]$. Then, we have $\mathbb{E}\left[w\left(\mathsf{DP}[G\oplus H,\varphi]\right)\right]\leq(1+\epsilon)\alpha\mathsf{opt}[G\oplus H,\varphi]$.
Lemma 42.
Let $0<\alpha\leq 1$, $\beta\in\mathbb{R}_{\geq 0}$ and assume
$\mathbb{E}\left[w(\mathsf{DP}[G\oplus H,\varphi])\right]$ $\leq\alpha\mathsf{opt}_{w}[G\oplus H,\varphi],$
$\mathbb{E}\left[(w+\delta\bm{1}_{u})(\mathsf{DP}[G\oplus H,\varphi])\right]$ $\leq\alpha\mathsf{opt}_{w+\delta\bm{1}_{u}}[G\oplus H,\varphi],$
$\mathsf{EM}\left(\mathsf{DP}_{w}[G,\varphi^{G}_{i}],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G,\varphi^{G}_{i}]\right)$ $\leq\beta$
holds for all $i\in[p]$. Then, we have
$\mathsf{EM}\left(\mathsf{DP}_{w}[G\oplus H,\varphi],\mathsf{DP}_{w+\delta\bm{1}_{u}}[G\oplus H,\varphi]\right)\leq 30\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.$
The following lemma corresponds to Lemma 34. The proof is the same as Lemma 40.
Lemma 43.
Let $\epsilon\in\mathbb{R}_{>0}$ and $h$ be the height of $T$. Our algorithm outputs a solution $X\in\mathsf{sat}(G,\varphi)$ such that $\mathbb{E}[w(X)]\leq(1+\epsilon)^{h}\mathsf{opt}[G,\varphi]$. The Lipschitz constant is bounded by $30(h+1)(1+\epsilon)^{h}\epsilon^{-1}\log(g(t,|\varphi|)\epsilon^{-1})$ for some computable function $g$.
Theorem 7 is also proved by the same proof as Theorem 2.